### v4

# Concentration Inequalities for Branching Random Walks

Abstract
While classical concentration inequalities are typically restricted to two special cases—independence and martingale difference sequences—we extend concentration inequalities to a much broader class of stochastic processes by relaxing these foundational conditions. Specifically, heuristically and in the language of calculus, while independence and the martingale difference property correspond to
$\frac{\partial y}{\partial t}=\text{constant},\quad\displaystyle\frac{\partial y}{\partial t}=0$
respectively, we relax these conditions to
a general setting that requires only the existence of a drift $\frac{\partial y}{\partial t}$ which is allowed to vary with the past state.
Furthermore, concentration inequalities are established for branching random walks.
Branching random walk,
keywords:
Keywords:
[…]
1 Introduction
1.1 Concentration inequalities
Let $(S_{i})_{i=1,2,...}$ be a real-valued random process, one-dimensional random walk. $S_{n}$ can be formulated as $S_{n}=X_{1}+X_{2}+...+X_{n}$; $X_{i}=S_{i}-S_{i-1}$. If some conditions are satisfied, among which independence is crucial, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT). For $S_{n}$ beyond the neighborhood, i.e., $|S_{n}-E(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [23]), also referred as Chernoff bound,
$Pr(|S_{n}-E(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
tells us how unlikely this occurrence is. Chernoff inequality is extended to the more general context of bounded martingale difference (say, $E(X_{n}\mid S_{n-1})=0$), referred as Azuma-Hoeffding inequality [2] (see [16] and [28] for surveys and references therein). From the point of view of our application and maybe many others’, Azuma-Hoeffding inequality and Chernoff inequality both give the same tight probability bound—square-exponential decay [^1] Thus the constant ‘$c$’ in the Chernoff bound is immaterial and the same symbol is used for simplicity., while the former has wider scope of applicability. After the classical results of Hoeffding and Azuma([23], [2]), refinements and extensions followed (e.g. [40], [28], [7], [8], [20], [25]). Known as concentration inequalities, these prove very useful and have a wide variety of applications in computer science, combinatorics, information theory (see e.g. [19], [4], [38], [29]).
However, for a long time, there had been no significant improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [40]. The reason might be two folds. First, there lacked motivation. Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results concealed the essential reason why the inequalities hold. In early 2000s [37] relaxed the boundedness to finite exponential moments; i.e. for a $\delta>0,E(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-E(S_{n})|\geq\lambda|)\leq e^{-\sim\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$ — not Chernoff bound. Later, motivated by the study of the free energy of a directed polymer in a random environment, [32] gave Chernoff bound under the hypothesis that, for a $\delta>0,E(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$). Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [25]. No attempt is known to get rid of martingale condition. The first major contribution of this paper is in this direction.
1.2 Notation and contribution
Let $U_{i}$ denote the partial sum of the first $i$ terms of $S_{n}=\sum_{j=1}^{n}X_{j}$, namely, $U_{i}=X_{1}+X_{2}+...+X_{i}$. We use lower case for variables scaled by $1/N$, where $N$ is a constant integer (in the context of $K$-SAT, NN represents the total number of variables). For instance, $u_{i}:=\sum_{j=1}^{i}X_{j}/N$, $x_{i}:=X_{i}/N$. For the sake of readability, in some places we use $Y_{n}$ for $S_{n}$.
We let $\mathbb{E}_{i}[\,\cdot\,]$ to abbreviate $\mathbb{E}[\,\cdot\mid u_{i}\,]$. Note that $\mathbb{E}_{i}[X_{t}]$ is a function of both $t$ and past state $(u_{i})_{i\leq t}$. To connect the calculus notation with the underlying stochastic process, we define
$\frac{\partial y}{\partial t}:=\frac{\mathbb{E}_{i}(y_{t})-\mathbb{E}_{i}(y_{t-1})}{1/N}=\mathbb{E}_{i}(Y_{t})-\mathbb{E}_{i}(Y_{t-1})=E_{i}(X_{t})$ (1.2)
where $Y_{t}=X_{1}+X_{2}+...+X_{t}$, and
$\frac{\partial^{2}y}{\partial u_{i}\,\partial t}:=\frac{\partial y(u_{i}^{\prime})/\partial t-\partial y(u_{i})/\partial t}{u_{i}^{\prime}-u_{i}}=\frac{\mathbb{E}(X_{t}\mid u_{i}^{\prime})-\mathbb{E}(X_{t}\mid u_{i})}{u_{i}^{\prime}-u_{i}}$ (1.3)
In the case of independence, $E_{i}(X_{t})=E(X_{t})$, we have
$\frac{\partial y}{\partial t}=\text{constant},$
In the case of martingale difference, $E_{i}(X_{t})=0$ for $i<t$ by induction and the law of iterated expectations,
$\frac{\partial y}{\partial t}=0$
Both conditions imply the mixed derivative vanishes:
$\frac{\partial^{2}y}{\partial u_{i}\,\partial t}=0$ (1.4)
While, traditionally, concentration inequalities require the condition (1.4), one of the primary contributions of this work is the relaxation of this classical constraint to the condition:
$\left|\frac{\partial^{2}y}{\partial u_{i}\,\partial t}\right|\leq L,\quad i=1,2,...,n$
i.e., $\frac{\partial^{2}y}{\partial u_{i}\,\partial t}$ is required only to be uniformly bounded, thereby allowing the drift $\frac{\partial y}{\partial t}$ to vary with $u_{i}$. This extends concentration results to general dependent processes with variable drift. For instance, the generation of random 3-SAT instances can be modelled as a branching random walk where, at each time step, a set of random clauses is applied to the current formula, producing $O(N^{3})$ progeny formulae from the current state; for details, see later in this paper. In this process, certain properties of the formula, such as the number of frozen variables, are history-dependent.
As a side benefit, this translation of classical probabilistic conditions from expectation-based to derivative-based formulations provides a calculus-based perspective on stochastic concentration phenomena.
1.3 Branching random walk (BRW)
In the literature, a branching random walk on the real line is described as follows (see, for example, [10] [13] [9]). In generation zero, an initial particle at the original place on the real line $\mathbf{R}$. It splits into a random number of child particles who form generation one. The children’s displacements, relative to their parent, correspond to a point process on $\mathbf{R}$. The children in turn split too to form the second generation, and so on. If the average split number (branching factor) is greater than one, with positive probability the number of the descendants grows exponentially through generations.
Current BRW studies typically address models where the offspring’s behavior is independent of that of their previous generation (e.g. [13], [9] [12], [24], [21]). The law of large numbers and central limit theorem type results about the distribution of position are established under certain conditions of independence (for instance, i.i.d of branching and walking). Yet unlike the situations of random walk, Chernoff bound is not known so far even in the case of i.i.d aforementioned, while the minimal (and maximal) is studied by many. Considering i.i.d offspring (and hence independent of the parent’s position), and iid displacement, Harris [22] conjectured that the distribution of the descendants’ position of the nth generation approaches Gaussian distribution. This conjecture was proved by [39], [33]. Its extended generation-dependent versions, where offspring and displacement distribution are determined by generation $n$, were proved by [11], [36], to mention a few. Problems where offspring and displacement are dependent of parents’ positions are studied in adhoc approaches, though (e.g. [46]). Concerning the deviation from the expectation, [9] establishes that for any $\delta>0$ the number of particles locating beyond $\mu$-$\delta$ is zero almost surely, $\mu$ denoting the scaled expectation. In another direction of estimating the rareness of the large deviation, extremum is well studied (e.g. [17] [27], [5], [6] [24], [1] [3] ). [17] showed tightness for $M_{n}-\mathbb{E}M_{n}$, where $M_{n}$ is minimum of nthnth generation. [27] gave probability bound for the deviation of $M_{n}$, i.e. $\Pr(M_{n}-Med_{n}>x)<e^{-\delta x}$, where $Med_{n}$ is the median of $M_{n}$. So far probability bound for only deviation $\epsilon$ of appropriate scale ($\epsilon>0$) is established. Little is known as to how sparse the population is at a distance of $O(\sqrt{n})$ far from the mean in terms of concentration inequalities.
1.4 Motivation and contribution
Concentration inequalities estimate the bound of the probability
$\int_{|S_{n}-E(S_{n})|>\lambda}p_{1}(x_{1}|u_{0})p_{2}(x_{2}|u_{1})...p_{n}(x_{n}|u_{n-1})dx_{1}dx_{2}...dx_{n}$ (1.5)
where $S_{n}=X_{1}+X_{2}+...+X_{n}$, $p_{1}(x_{1})p_{2}(x_{2})...p_{n}(x_{n})$ is the probability density of $(x_{1},...x_{n})$ (here and throughout we use $p_{i}(x_{i})$ for $p_{i}(x_{i}|u_{i-1})$ for notational simplicity).
In algorithm analysis and random graphs study, which are described as graph evolution stochastic process (see e.g. [34], [42], [43], [44], [45] and references therein), the differential equation method (DEM) started in 80’s. Recently DEM has obtained some important results and received increasingly attention. The procedure of DEM consists of (a) verifying that the random parameters concerned concentrate around their means, and (b) finding solutions to the differential equations satisfied the means. Concentration inequalities (e.g. Heoffding-Azuma inequalities) are used to prove concentration phenomena in DEM.
However, concentration inequalities can only deal with random walk process (RW), while there are large class of problems where the random process is branching random walk (BRW), for example K-SAT/q-COL to mention a few ([31]). The probability density of BRW is ([30] [31])
$\int_{\left|\,S_{n}-E(S_{n})\,\right|>\lambda}\frac{m_{1}(u_{0})p_{1}(x_{1})...m_{n}(u_{n-1})p_{n}(x_{n})}{\displaystyle\int m_{1}(u_{0})p_{1}(x_{1})...m_{n}(u_{n-1})p_{n}(x_{n})dx_{1}...dx_{n}}\,dx_{1}...dx_{n}$ (1.6)
where $m_{i}$ is branching factor. Except for special $m_{i}$ (e.g., $m_{i}$ is independent of $u_{i-1}$; e.g. $m_{i}=1$), (1.6) can not be reduced to (1.5); thus, current concentration inequalities are not applicable. In this paper we present concentration inequalities in the context of BRW—the second major contribution of this paper. Although originally motivated by the project to attack the problems of K-SAT/q-COL threshold location, they may be of independent interest. For the applications, brief list of the results obtained is presented in this paper; for details, the readers are referred to [31], a companion paper.
2 An extension of Heoffding-Azuma inequality
For warming up, let’s present a “downgraded” proof for Azuma’s inequality. Possibly this proof could be found in the Azuma’s trash can. Let’s first introduce a simple and elegant inequality due to [32] and [41]. Given here is the version from [41] which improves [32].
Lemma 2.1.
If $\mathbb{E}(X)=0$, then for all $\delta>0$ and $|t|/\delta\in[0,1]$
$\mathbb{E}\left[e^{tX}\right]\leq 1+(t/\delta)^{2}\,\mathbb{E}\left[e^{\delta|X|}\right]\leq\exp\left(t^{2}/\delta^{2}\cdot\mathbb{E}[\,e^{\delta|X|}\,]\right)$
In particular, when $\delta=1$,
$\mathbb{E}\left[e^{tX}\right]\leq\exp{\big(t^{2}\mathbb{E}[e^{|X|}]\big)}$ (2.1)
Proof.
$\mathbb{E}(e^{tX})$ $=1+\mathbb{E}[tX]+\mathbb{E}\left[\frac{t^{2}X^{2}}{2!}+\frac{t^{3}X^{3}}{3!}+\cdots\right]$
$=1+\mathbb{E}\left[\frac{(t/\delta)^{2}\cdot(\delta X)^{2}}{2!}+\frac{(t/\delta)^{3}\cdot(\delta X)^{3}}{3!}+\cdots\ \right]$
$\leq 1+(t/\delta)^{2}\mathbb{E}\left[\frac{|\delta X|^{2}}{2!}+\frac{|\delta X|^{3}}{3!}+\cdots\ \right]\quad\text{(employing $|\,t\,|/\delta\leq 1$})$
$\leq 1+(t/\delta)^{2}\mathbb{E}\left[e^{\delta|X|}\right]\leq\exp\left(t^{2}/\delta^{2}\cdot\mathbb{E}[\,e^{\delta|X|}\,]\right)$
The proof is concluded. ∎
Theorem 2.1.
(Azuma’s Inequality) Let $\{X_{i}\}_{i=1}^{n}$ be a sequence of martingale difference, with $|X_{i}|\leq 1$ (without loss of generality), and $S_{n}=X_{1}+X_{2}+...+X_{n}$, then
$Pr(|S_{n}-ES_{n}|\geq\lambda)\leq 2\exp\Big(-\frac{\lambda^{2}}{2nc^{2}}\Big)$
To prove the inequality, Azuma ([2]) employed the following inequalities, for $|X|\leq c$,
$e^{tX}\leq\frac{e^{tc}-e^{-tc}}{2c}X+\frac{e^{tc}+e^{-tc}}{2}\quad\text{(due to convexity of $e^{tX}$)}$ (2.2)
and
$\frac{e^{tc}+e^{-tc}}{2}\leq e^{t^{2}c^{2}/2}\quad\text{($\cosh(x)<e^{x^{2}/2}$)}$ (2.3)
to get sub-gaussian condition
$E(e^{tX})\leq e^{t^{2}c^{2}/2}~~~~~~\mbox{for}~\mathbb{E}(X)=0$
as adopted by textbooks. Sub-gaussian condition is a key ingredient in the proof of concentration inequalities, as this ultimately leads to a bound of the form
$\exp({\sim nt^{2}-t\lambda})$
which is exponentially small under the choice $t\sim n^{-1/2}$, even for $\lambda\ll n$.
Proof (Downgraded version of Azuma’s inequality).
We only consider $\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$, as the proof of the reverse $\mathbb{P}(S_{n}-\mathbb{E}S_{n}\leq-\lambda)$ is exactly the same. By Markov’s inequality, inductive application of Lemma 2.1 and optimization of $t$, we have
$\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$ $\leq e^{-t\lambda}\cdot E\{e^{tS_{n-1}}\cdot\mathbb{E}(e^{tX_{n}}\mid X_{1},...,X_{n-1})\}$
[…]
$\leq e^{-t\lambda}\cdot\mathbb{E}\{e^{tS_{n-1}}\cdot e^{t^{2}e^{c}}\}~~~~~~~\mbox{(}|X_{i}|\leq c\mbox{)}$
$\leq\cdots\leq\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}\quad\mbox{(cf. Azuma's $\exp{\left(\displaystyle-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$)}$
The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [32] and [41]) without compromising strength of results; from the practical point of view, the difference between $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ and $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$ is immaterial – both are square-exponential decay. In addition, it is demonstrated that nifty techniques are not necessary for obtaining concentration inequalities.
Next, we will show what else can make the concentration inequality hold, other than independence and martingale difference. To this end, define Doob’s (or McDiarmid’s) martingale
$d_{i}$ $=\mathbb{E}_{i}(X_{i}+\cdots+X_{n})-\mathbb{E}_{i-1}(X_{i}+\cdots+X_{n})$ (2.4)
It is easy to check
$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\quad$
$E_{i-1}(d_{i})=0$. Notice
$d_{i}$ $=X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)$ (2.5)
Given $u_{i-1}$, $E_{i}(X_{l})$ is a function of $X_{i}$. $E_{i-1}(X_{l})=E_{X_{i}}\{E_{i}(X_{l})\}$ where $E_{X_{i}}(\cdot)$ takes expectation with respect to $X_{i}$. We have
Lemma 2.2.
If
$\left|\,\mathbb{E}_{i-1}(X_{\ell}\mid X_{i})-\mathbb{E}_{i-1}(X_{\ell}\mid X_{i}^{\prime})\,\right|\leq L\displaystyle\frac{|X_{i}-X_{i}^{\prime}|}{N},\quad\text{for $\ell>i$}\;$ (2.6)
then
$|\mathbb{E}_{i}(X_{\ell})-\mathbb{E}_{i-1}(X_{\ell})|\leq L\frac{|X_{i}|}{N}+L\frac{\mathbb{E}_{i-1}(\,|X_{i}|\,)}{N}$ (2.7)
Note, $\mathbb{E}_{i}(X_{\ell})=\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
Proof (Lemma 2.2).
Let
$\mathbb{E}_{X_{i}^{\prime}}(\cdot)\equiv\mathbb{E}_{X_{i}^{\prime}}(\,\cdot\mid u_{i-1}),\quad X_{i}^{\prime}\sim\mathcal{L}(X_{i}\mid u_{i-1}).$
With this notation,
$\bigl|\mathbb{E}_{i}\left(X_{\ell}\right)-\mathbb{E}_{i-1}\left(X_{\ell}\right)\bigr|$ $=\Bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{X_{i}^{\prime}}\!\left[\,\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\right]\Bigl|$
$=\Bigl|\,\mathbb{E}_{X_{i}^{\prime}}\big[\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\big]\,\Bigl|$
$\leq\mathbb{E}_{X_{i}^{\prime}}\Big(\,\bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\bigl|\,\Big)$
$\leq\mathbb{E}_{X_{i}^{\prime}}\!\left[\,L\frac{|X_{i}-X_{i}^{\prime}|}{N}\,\right],\quad\text{$\mathbb{E}_{i-1}\left[\,X_{\ell}\mid X_{i}\,\right]$ is L-Lipschitz in $X_{i}$}$
$\leq\frac{L}{N}|X_{i}|+\frac{L}{N}E_{X_{i}^{\prime}}\left(\,|X_{i}^{\prime}|\,\right)$
$=L\frac{|X_{i}|}{N}+L\frac{\mathbb{E}_{i-1}\left(\,|X_{i}|\,\right)}{N}$
The claim is true. ∎
Lemma 2.3.
Let $A$ denote $1+L\frac{M}{N}$, where $M=O(N)$. If $\mathbb{E}_{i-1}\left[\,X_{\ell}\mid X_{i}\,\right]$ is L-Lipschitz in $X_{i}$, namely
$\left|\,\mathbb{E}_{i-1}(X_{\ell}\mid X_{i})-\mathbb{E}_{i-1}(X_{\ell}\mid X_{i}^{\prime})\,\right|\leq L\displaystyle\frac{|X_{i}-X_{i}^{\prime}|}{N},\quad\text{for $\ell>i$}$ (2.8)
and $\mathbb{E}[\,e^{\delta A|X_{i}|}\mid u_{i-1}\,]\leq K$ (exponential moment existence) for a constant $\delta>0$ , $i=1,2,$..., then for $n\leq M$
$\Pr\big(\,|\,S_{n}-\mathbb{E}S_{n}\,|\geq\lambda\,\big)\leq 2\exp{\Big(\!-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\Big)}$ (2.9)
Obviously, Poisson distribution $\text{Poisson}(\lambda_{i})$ meets the exponential moment condition.
Proof of Lemma 2.3.
By definition (2.5),
$\mathbb{E}_{i-1}\exp\left(\delta|d_{i}|\right)=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
$\leq\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|\mathbb{E}_{i-1}(X_{i})|+\delta\sum_{\ell=i+1}^{n}\big|\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big|\Big)\Big]$
[…]
$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|X_{i}|L\frac{M}{N}+\delta\,\mathbb{E}_{i-1}(\,|X_{i}|\,)\cdot L\frac{M}{N}\Big)\Big]$
$=\exp\left(\mathbb{E}_{i-1}\delta|X_{i}|\,+\,\delta L\frac{M}{N}\mathbb{E}_{i-1}|X_{i}|\right)\cdot\mathbb{E}_{i-1}\left[\exp\left(\delta|X_{i}|\Big(1+L\frac{M}{N}\Big)\right)\right]$
$=\exp\big(\mathbb{E}_{i-1}\delta\cdot A|X_{i}|\big)\cdot\mathbb{E}_{i-1}(e^{\delta A|X_{i}|})\leq K^{2}~~~~\mbox{(Jensen's inequality and the hypothesis)}$
Applying the exponential Markov inequality $\Pr[X\geq\lambda]\leq e^{-t\lambda}E(e^{tX})$ and noticing
$E_{i-1}(d_{i})=0$
we have
$\Pr\big(S_{n}-\mathbb{E}S_{n}\geq\lambda\,\big)$
$\leq e^{-t\lambda}\,\mathbb{E}\!\left[e^{\,t(S_{n}-\mathbb{E}S_{n})}\right]=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(d_{1}+d_{2}+\cdots+d_{n})}\right]=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(d_{1}+\cdots+d_{n-1})}\;\mathbb{E}_{n-1}\left(e^{td_{n}}\right)\right]$
[…]
$\leq e^{-t\lambda}\,e^{\,t^{2}K^{2}/\delta^{2}}\;\mathbb{E}\!\left[e^{t(d_{1}+\cdots+d_{n-2})}\;\mathbb{E}_{n-2}\!\left(e^{td_{n-1}}\right)\right]$
$\quad\vdots$
$\leq e^{-t\lambda}\,e^{\,nt^{2}K^{2}/\delta^{2}}.$
Optimizing in $t$, we obtain
$\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)\leq\exp\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right).$
For the other half side of the concentration inequality, we refer the fact that Lemma 2.1 holds for negative $t$ as well. The proof is exactly the same as above, which we omit. ∎
Remarks.
  • Condition (2.8) in Lemma 2.3 is equivalent to the following Lipschitz continuity
    $\left|\mathbb{E}(X_{\ell}\mid u_{i})-\mathbb{E}(X_{\ell}\mid u_{i}^{\prime})\right|\leq L\left|u_{i}-u_{i}^{\prime}\right|\quad\mbox{for $i<l$}$ (2.10)
    where
    $u_{i}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}),\quad u_{i}^{\prime}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}^{\prime}).$
  • The condition of Lipschitz continuity (2.10) may be rephrased “uniqueness of solution”; i.e. small difference in initial position can not yield big difference in the trajectory.
  • In the case of independence, $\mathbb{E}_{i}[X_{\ell}]=\mathbb{E}[X_{\ell}]$, and hence $\mathbb{E}_{i}[X_{\ell}]-\mathbb{E}_{i-1}[X_{\ell}]$ = 0. In the case of martingale difference sequence, $\mathbb{E}_{i}[X_{\ell}]=0$ for $\ell>i$. In both cases,
    $\sum_{\ell>i}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)=0$
    and therefore $d_{i}$ (see, (2.5)) is reduced to $X_{i}-\mathbb{E}X_{i}$ (independence) and to $X_{i}$ (martingale difference), respectively. In these cases, the condition of bounded difference or the existence of exponential moment alone suffices to imply the concentration inequalities.
  • The condition of Lipschitz continuity (2.10) may be rephrased “uniqueness of solution”; i.e. small difference in initial position can not yield big difference in the trajectory.
  • Recall (1.3), the notations defined in the section of Introduction. In the notation of calculus, the condition (2.10) is
    $\left|\frac{\partial}{\partial u_{i}}\left(\frac{\partial y}{\partial t}\right)\right|\leq L.$ (2.11)
  • Two special cases of (2.10) are independence where $\frac{\partial y}{\partial t}$ is constant and martingale difference where $\frac{\partial y}{\partial t}=0$.
The following theorem provides easily verifiable conditions for concentration inequality in general cases where
$\frac{\partial}{\partial u_{i}}\left(\frac{\partial y}{\partial t}\right)\neq 0$
i.e. $\frac{\partial y}{\partial t}$ is not constant with $u_{i}$ varying.
$\mathbb{E}(X_{i+1}|u_{i})$ is a function of $u_{i}$. Let $\mu_{i}(u_{i})$ denote $E(X_{i+1}|u_{i})$ and $\nu_{i}(u_{i})$ denote $E(X_{i+1}^{2}|u_{i})$;
they are functions of $u_{i}$.
Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on. We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the region in question is distant enough from singularities).
Theorem 2.2.
If $E\big(e^{\delta A|X_{i}|}\mid u_{i-1}\big)\leq K$ (existence of an exponential moment), then in the region where $\mu_{i}$ and $\nu_{i}$ are smooth, the concentration inequality holds.
In the notation of partial derivatives, the condition in Theorem 2.2 states that
$\frac{\partial\mu(t,u)}{\partial u},\quad\frac{\partial^{2}\mu(t,u)}{\partial u^{2}},\quad\frac{\partial\nu(t,u)}{\partial u},\cdots$
are uniformly bounded in the region of smoothness (i.e., away from singularities), in addition to the exponential moment condition. These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE). From the PDE we have every partial derivatives of $\bar{u}$, and $u$ is close to $\bar{u}$ with exponentially rare and negligible exceptions. Note, the condition (2.10) says that small changes in distant past generations won’t lead to violently change in the current step size. Theorem 2.2 tells us, in terms of PDE of $\bar{u}$, what local behavior of $u$ can meet (2.10), for the concentration inequality to hold.
Proof.
We just need to prove Lipschitz continuity (2.10). The proof is proceeded by induction. The base case holds, i.e. $|E(X_{i+1}|u_{i})-E(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
In the following, we investigate stochastic process of $u_{n}$, $n>i$. With $u_{i}$ is given, which may be treated as a deterministic variable, we write $\mathbb{E}$ for $\mathbb{E}_{i}$ .
$\mathbb{E}(X_{n+1}\mid u_{n})$ is a function of $u_{n}$, and $u_{n}$ itself is a function of $u_{i}$. Therefore $\mathbb{E}(X_{n+1}|u_{n})$ is ultimately a function of $u_{i}$, which we denote by $\mu_{n}$. Taking expectation of the Taylor series expansion of $\mu_{n}(u_{n})$ around the mean of $u_{n}$, gives
$\mathbb{E}(X_{n+1})\,$ $=\mathbb{E}\left[\mathbb{E}(X_{n+1}\mid u_{n})\right]=\,\mathbb{E}[\mu_{n}(u_{n})]$
$=\mu_{n}(\bar{u}_{n})~+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathbb{E}(u_{n}-\bar{u}_{n})^{2}+\frac{1}{3!}\mathbb{E}\left[\mu_{n}^{\prime\prime\prime}(\xi)(u_{n}-\bar{u}_{n})^{3}\right]$
By induction, $u_{n}$ exhibits concentration. It follows
[…]
By the hypothesis of “smoothness”, $\mu_{n}^{\prime\prime\prime}(\xi)$ exists. Hence
$\big|\mathbb{E}\left(\mu_{n}^{\prime\prime\prime}(\xi)(u_{n}-\bar{u}_{n})^{3}\right)\big|\lesssim\;{N^{-3/2}}$
and therefore we have
$\mathbb{E}(X_{n+1})=\mu_{n}(\bar{u}_{n})~+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\frac{C}{N^{3/2}}$ (2.12)
where $\mathrm{Var}_{n}$ denotes $\mathbb{E}(u_{n}-\bar{u}_{n})^{2}$. Similarly, for $\nu$ we have
$\mathbb{E}(X_{n+1}^{2})=$ $\nu_{n}(\bar{u}_{n})+\frac{1}{2}\nu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+~\frac{C}{N^{3/2}}$ (2.13)
From
$\mathrm{Var}_{n+1}=\mathbb{E}\big(u+x_{i+1}\cdots+x_{n}-(u+\bar{x}_{i+1}+\cdots+\bar{x}_{n})+(x_{n+1}-\bar{x}_{n+1})\big)^{2},$
the following recurrence equation about $\mathrm{Var}_{n+1}$ holds.
$\mathrm{Var}_{n+1}=\mathrm{Var}_{n}+2\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]+\frac{1}{N^{2}}\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ (2.14)
We first estimate the second term in the right hand-side of (2.14).
$\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mathbb{E}[X_{n+1}(u_{n}-\bar{u}_{n})]-\bar{x}_{n+1}(\bar{u}_{n}-\bar{u}_{n})$
$=$ $\frac{1}{N}\mathbb{E}[\mu_{n}(u_{n})(u_{n}-\bar{u}_{n})]$
Again applying Taylor expansion to $\mu_{n}(u_{n})$, i.e.
$\mu_{n}(u_{n})=\mu_{n}(\bar{u}_{n})+\mu_{n}^{\prime}(\bar{u}_{n})(u_{n}-\bar{u}_{n})+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})(u_{n}-\bar{u}_{n})^{2}+\sim(u_{n}-\bar{u}_{n})^{3}$
we have, with the term of $\sim\mathbb{E}(u_{n}-\bar{u}_{n})^{4}$ neglected,
$\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mu_{n}(\bar{u}_{n})\mathbb{E}(u_{n}-\bar{u}_{n})+\frac{1}{N}\mu_{n}^{\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\frac{1}{2N}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$
$=$ $\frac{1}{N}\mu_{n}^{\prime}(\bar{u}_{n})\mathrm{Var}_{n}$
where the term of $\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$ is neglected.
Now let’s estimate $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$. From (2.13) and (2.12), we obtain
$\mathbb{E}(X_{n+1}^{2})-\overline{X}_{n+1}^{2}=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-[\mu_{{n}}(\bar{u}_{n})+\frac{1}{2}\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}]^{2}$
$=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-\mu_{{n}}^{2}(\bar{u}_{n})-\mu_{{n}}(\bar{u}_{n})\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}$
$=$ $\nu_{{n}}(\bar{u}_{n})-\mu_{{n}}^{2}(\bar{u}_{n})+\left(\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})-\mu_{n}(\bar{u}_{n})\mu_{n}^{\prime\prime}(\bar{u}_{n})\right)\mathrm{Var}_{n}$
For illustration, at $i+1$
$\mathbb{E}(X_{i+1}-\overline{X}_{i+1})^{2}=\mathbb{E}(X_{i+1}^{2})-\mu_{i}^{2}=\nu_{i}(u_{i})-\mu_{i}^{2}(u_{i})=\mathrm{Var}(X_{i+1})$
where $\mathrm{Var}(X_{i})=0$ because $u_{i}$ is treated as a deterministic variable.
By substituting $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ and $\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]$ into (2.14), we have a recurrence relation for $n\geq i$,
$\mathrm{Var}_{n+1}=$ $\mathrm{Var}_{n}+\frac{2}{N}\mu_{n}^{\prime}(\bar{u}_{n})\mathrm{Var}_{n}~+\frac{1}{N^{2}}\left[\nu_{{n}}(\bar{u}_{n})-\mu_{{n}}^{2}(\bar{u}_{n})+\left(\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})-\mu_{n}(\bar{u}_{n})\mu_{n}^{\prime\prime}(\bar{u}_{n})\right)\mathrm{Var}(u_{n})\right]$
$=$ $\mathrm{Var}_{n}+\frac{\mathrm{Var}_{n}}{N}\left(2\mu_{n}^{\prime}(\bar{u}_{n})~+\frac{1}{2N}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})-\frac{1}{N}\mu_{n}(\bar{u}_{n})\mu_{n}^{\prime\prime}(\bar{u}_{n})\right)+\frac{1}{N^{2}}\left(\nu_{n}(\bar{u}_{n})-\mu_{n}^{2}(\bar{u}_{n})\right)$
and $\mathrm{Var}_{i}=0$. With the argument $\bar{u}_{n}$” omitted for brevity:
$\mathrm{Var}_{n+1}=$ $\mathrm{Var}_{n}\left[1+\frac{1}{N}\left(2\mu_{n}^{\prime}+\frac{1}{2N}\nu_{n}^{\prime\prime}-\frac{1}{N}\mu_{n}\mu_{n}^{\prime\prime}\right)\right]+\frac{1}{N^{2}}(\nu_{n}-\mu_{n}^{2})$
Define
$a_{n}=\displaystyle 1+\frac{1}{N}\left(2\mu_{n}^{\prime}+\frac{1}{2N}\nu_{n}^{\prime\prime}-\frac{1}{N}\mu_{n}\mu_{n}^{\prime\prime}\right),\qquad b_{n}=\frac{\nu_{n}-\mu_{n}^{2}}{N^{2}}$
Then, the above recursion for $\mathrm{Var}$ gives
$\mathrm{Var}_{n+1}=$ $\sum_{j=i+1}^{n-1}b_{j}\cdot a_{j+1}...a_{n}+b_{n}=\sum_{j=i+1}^{n-1}b_{j}\cdot a_{j+1}...a_{n}+\frac{1}{N^{2}}(\nu_{n}-\mu_{n}^{2})$
$=$ $\frac{1}{N^{2}}\left[\sum_{j=i+1}^{n-1}(\nu_{j}-\mu_{j}^{2})\prod_{k=j+1}^{n}\left(1+\frac{1}{N}\Big(2\mu_{k}^{\prime}+\frac{1}{2N}\nu_{k}^{\prime\prime}-\frac{1}{N}\mu_{k}\mu_{k}^{\prime\prime}\Big)\right)+\nu_{n}-\mu_{n}^{2}\right]$
$=$ $\frac{1}{N^{2}}\left[\sum_{j=i+1}^{n-1}\left(\nu_{j}-\mu_{j}^{2}\right)\prod_{k=j+1}^{n}\left(1+\frac{2\mu_{k}^{\prime}}{N}\right)\right]$
(higher order terms are neglected)
Given above, it is easy to check that, in the region where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$. Further, by induction, $\frac{d\mathrm{Var}_{n+1}}{du}$ exists and its order is $O(1/N)$.
From the base that $\Delta u_{i}=u_{i}-u_{i}^{\prime}$, and the inductive hypothesis that $\Delta\bar{u}_{n-1}$ is small, it follows from (2.12) that
$\Delta\mathbb{E}(X_{n})$ $=\Delta\mu_{n-1}~+~\Delta\frac{1}{2}\mu_{n-1}^{\prime\prime}\mathrm{Var}_{n-1}~+~\frac{C}{N^{3/2}}$
$=\mu_{n-1}^{\prime}\Delta\bar{u}_{n-1}+\frac{1}{2}\mu_{n-1}^{\prime\prime\prime}\mathrm{Var}_{n-1}\Delta\bar{u}_{n-1}+\frac{1}{2}\mu_{n-1}^{\prime\prime}\frac{d\mathrm{Var}_{n-1}}{du}\Delta\bar{u}_{n-1}~+~O(\Delta\bar{u}_{n-1}^{2})$
$=\mu_{n-1}^{\prime}\Delta\bar{u}_{n-1}+\frac{1}{2}\mu_{n-1}^{\prime\prime\prime}O\Big(\frac{1}{N}\Big)\Delta\bar{u}_{n-1}+\frac{1}{2}\mu_{n-1}^{\prime\prime}O\Big(\frac{1}{N}\Big)\Delta\bar{u}_{n-1}~+~O(\Delta\bar{u}_{n-1}^{2})$
$=\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O\left(N^{-1}\right)+O(\Delta\bar{u}_{n-1})\Big)$
where the 3rd term in (2.12), of order $N^{-3/2}$, is omitted since it yields $\sim\displaystyle O(N^{-3/2})\Delta\bar{u}_{n-1}$ in the above equation. From $u_{n}=u_{n-1}+X_{n}/N$ and then
$\bar{u}_{n}=\bar{u}_{n-1}+\frac{1}{N}E(X_{n})$
we have
$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta E(X_{n})=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O\left(N^{-1}\right)+O(\Delta\bar{u}_{n-1})\Big)$
Thus,
$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}\left(1+\frac{1}{N}\mu_{n-1}^{\prime}+o\Big(\frac{1}{N}\Big)\right)$
Iterating on $n$ yields
$|\,\mathbb{E}(X_{n}\mid u_{i})-\mathbb{E}(X_{n}\mid u_{i}^{\prime})\,|\leq L|u_{i}-u_{i}^{\prime}|\quad\text{(where $u_{i}-u_{i}^{\prime}=\Delta u_{i}$)}$
This concludes the proof. ∎
We see from the above proof that Lipschitz continuity of $\mathbb{E}(f(u_{n}))$ is ultimately down to the existence of the first several order derivatives of $f$.
3 Concentration inequalities in the context of BRW
Let’s first fix some notations. It is suggested [30] that BRW be defined by a sequence of pairs
$\Big(m_{i}(u_{i-1}),p(X_{i}\mid u_{i-1})\Big)_{i=1,2,\cdots,n}$
where the branching factor $m_{i}$ (also called birth-rate) is the expectation of offsprings which is dependents on the birthplace $u_{i}$, and
$p(X_{i}\mid u_{i-1})$
is the probability density function (pdf) of children’s displacement which is dependent of the birthplace as well. Here $m_{i}$ can be either greater than 1 (so called supercritical) or less than 1 (subcritical). If $m_{i}$ is constant, the BWR is equivalent to random walk process where $m_{i}=1$. In other words, random walk and branching random walk is mathematically equivalent if $m_{i}$ is constant. Probability density function (pdf) of BRW can be found in (1.6). In K-SAT, for instance, $m_{n}=1-u_{n}^{\!{}_{K}}$. That is, if a particle’s position is $u_{n}$ the size of its offsprings is $1-u_{n}^{\!{}_{K}}$.
Intuitively, if the branching is “squeezing” — i.e. if the branching factor is smaller in the regions farther from the mean path — then the BRW is more concentrated than without branching (or branching factor = 1). In other words, if you squeeze the population towards its mean by reducing the birth rate in the remote regions, then the population distribution becomes more concentrated. In context of K-SAT, the instances with less-constraint have more descendants than those with more constraints; at least, the reverse cannot be true.
Formally, let $Z(u_{i})$ be the total progeny of the particle at position $u_{i}$. Given fixed $u_{i-1}$, this is a function of $X_{i}$, which we write as $Z(X_{i})$. We make the following assumption.
Assumption (negative association). If $f(\cdot)$ is a monotonically increasing function, then
$\int f(|X_{i}|)\,Z(X_{i})\,P(dX_{i})\;\leq\;\left(\int f(|X_{i}|)\,P(dX_{i})\right)\left(\int Z(X_{i})\,P(dX_{i})\right).$ (3.1)
If $Z(X_{i})$ is decreasing in $X_{i}$, then the negative association inequality (3.1) holds.
In a random walk (RW), a particle moves endlessly — it never dies, never mates, and therefore never reproduces. The expectation of $(\cdot)$ is
$\mathbb{E}[\,\cdot\,]=\int(\,\cdot\,)P(dX_{1}\cdots dX_{n})$
where $P(dX_{1}\cdots dX_{n})=p_{1}(X_{1})\cdots p_{n}(X_{n})\ dX_{1}\cdots dX_{n}$. In BRW, each particle reproduces descendants generation by generation. The average $(\cdot)$ over the whole population is
$\mathbb{E}^{\!{}^{(M)}}[\,\cdot\,]=\frac{\displaystyle\int(\,\cdot\,)\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}{\displaystyle\int\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}$
where $m_{i}=m_{i}(u_{i-1}+X_{i}/N)$. This is dependent of the future generations, whereas statistics in RW is independent of future. Therefore any statistical measure in BRW is generation dependent.
Define conditional expectation,
$\mathbb{E}{\!{}^{(M)}}[\,\cdot\mid u_{i}\,]:=\mathbb{E}_{i}^{\!{}^{(M)}}[\,\cdot\,]=\frac{\displaystyle\int(\,\cdot\,)\prod_{j=i+1}^{M}m_{j}\ P(dX_{i+1}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i})}{\displaystyle\int\prod_{j=i+1}^{M}m_{j}\ P(dX_{i+1}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i})}$ (3.2)
For example,
$\begin{split}&\mathbb{E}^{\,\!{}^{(M)}}_{i}[\,X_{\ell}\,]:=\mathbb{E}{\!{}^{(M)}}[\,X_{\ell}\mid X_{1},...,X_{i}\,]=\frac{\displaystyle\int X_{\ell}\prod_{j=i+1}^{M}m_{j}\ P(dX_{i+1}\cdots dX_{\!{}_{M}})}{\displaystyle\int\prod_{j=i+1}^{M}m_{j}\ P(dX_{i+1}\cdots dX_{\!{}_{M}})}\end{split}$ (3.3)
where $u_{i}=u_{0}+X_{1}/N+\cdots+X_{i}/N$.
Define
$P^{\,\!{}^{(M)}}_{i}(dX_{i+1}\cdots dX_{M})=\frac{m_{i+1}(u_{i})\cdots m_{\!{}_{M}}(u_{\!{}_{M-1}})\ P(dX_{i+1}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i})}{\displaystyle\int m_{i+1}(u_{i})\cdots m_{\!{}_{M}}(u_{\!{}_{M-1}})\ P(dX_{i+1}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i})}$
So the equality (3.3) can be written as
$\mathbb{E}{\!{}^{(M)}}[\,X_{\ell}\mid X_{1},...X_{i}\,]=\int X_{\ell}\ P^{\!{}^{(M)}}_{i}(dX_{i+1}\cdots dX_{\ell}\cdots dX_{M})$ (3.4)
When the branching factors $m_{j}$ are constant, the BRW conditional measure reduces to the underlying RW conditional measure:
$P^{\!{}^{(M)}}_{i}(dX_{i+1}\cdots dX_{M})={\ P(dX_{i+1}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i})}$
and equality (3.4) becomes
$\mathbb{E}{\!{}^{(M)}}[\,X_{\ell}\mid X_{1},...X_{i}\,$] $=\int X_{\ell}\ P(dX_{i+1}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i})$
$=\int X_{\ell}\ P(dX_{i+1}\cdots dX_{\ell}\mid X_{1},...,X_{i})$
The second equality holds because, when the branching factors are constant, the future process does not affect the earlier behaviour.
By the law of iterated expectations
$\mathbb{E}^{\,\!{}^{(M)}}[\,\cdot\,]=\mathbb{E}^{\,\!{}^{(M)}}\big(\,\mathbb{E}_{i}^{\,\!{}^{(M)}}[\,\cdot\,]\,\big)$
that is, the BRW expectation satisfies the usual law of total expectation, despite being defined through weighted (branching) measures. We distinguish between two types of expectation. One is the standard (unweighted) expectation
$\mathbb{E}[\,\cdot\,\mid u_{i}\,]:=\int(\,\cdot\,)\,P(dX_{i+1})=\int(\,\cdot\,)\,p_{i+1}(X_{i+1}\mid u_{i})dX_{i+1}$
and the other is the BRW-weighted expectation $E^{\,\!{}^{(M)}}$ or $E^{\,\!{}^{(M)}}[\,\cdot\mid u_{i}\,]$
Define
$m_{i}(u_{i-1}):=\mathbb{E}[\,m_{i}(u)\mid u_{i-1}\,]$
where $u_{i+1}=u_{i}+X_{i+1}/N$.
The following lemma is the BRW counterpart of Lemma 2.3. The proof thereof is exactly the same as Lemma 2.3, which we omit.
Lemma 3.1.
Let $A$ denote $1+L\displaystyle\frac{M}{N}$, where $M=O(N)$. If
$\left|\,\mathbb{E}^{\,\!{}^{(M)}}_{i-1}(X_{\ell}\mid X_{i})-\mathbb{E}^{\,\!{}^{(M)}}_{i-1}(X_{\ell}\mid X_{i}^{\prime})\,\right|\leq L\displaystyle\frac{|X_{i}-X_{i}^{\prime}|}{N},\quad\text{for $\ell>i$}$ (3.5)
and
$\mathbb{E}^{\,\!{}^{(M)}}_{i-1}(e^{\delta A|X_{i}|})\leq K\quad\textnormal{(exponential moment exists)}$ (3.6)
for a constant $\delta>0$ , $i=1,2,\cdots,$ then
$Pr[\ |S_{n}-\mathbb{E}^{{(M)}}(S_{n})\mid\geq\lambda]\leq 2\exp{\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)},\quad\text{$n\leq M$ }$ (3.7)
Remark.
  • Condition (3.5) in Lemma 3.1 is equivalent to the following Lipschitz continuity
    $\left|\mathbb{E}^{\,\!{}^{(M)}}(X_{\ell}\mid u_{i})-\mathbb{E}^{\,\!{}^{(M)}}(X_{\ell}\mid u_{i}^{\prime})\right|\leq L\left|u_{i}-u_{i}^{\prime}\right|\quad\mbox{for $i<l$}$
    where
    $u_{i}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}),\quad u_{i}^{\prime}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}^{\prime}).$
The following theorem is the counterpart of Theorem 2.2, which gives easily verifiable conditions for BRW concentration inequality (3.7).
Theorem 3.1.
If $E(e^{\delta A|X_{i}|}\mid u_{i-1})\leq K$, then in the area where $m_{i}(u)$, $\mu_{i}(u)$ and $\nu_{i}(u)$ are smooth, the concentration inequality holds.
Proof.
Let $Z_{i}(u_{i})$ denote the population descended from the birthplace $u_{i}$, i.e.
$Z_{i}(u_{i})=\int m_{i+1}...m_{\!{}_{M}}P(dX_{i+1}...dX_{\!{}_{M}}\mid u_{i}),\quad u_{i}=u_{i-1}+\frac{X_{i}}{N}$
From the definition of $E^{(M)}_{i-1}$$\cdot\,$] (see (3.2)),
$\mathbb{E}_{i-1}^{\!{}^{(M)}}[\,e^{\delta A|X_{i}|}\,]$ $=\frac{\displaystyle\int e^{\delta A|X_{i}|}\,m_{i}\cdots m_{\!{}_{M}}\ P(dX_{i}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i-1})}{\displaystyle\int m_{i}\cdots m_{\!{}_{M}}\ P(dX_{i}\cdots dX_{\!{}_{M}}\mid X_{1},...,X_{i-1})}$
$=\frac{\displaystyle\int e^{\delta A|X_{i}|}\,m_{i}(u_{i-1}+X_{i})Z_{i}\ P(dX_{i}|u_{i-1})\;}{\displaystyle\int m_{i}(u_{i-1}+X_{i})Z_{i}\ P(dX_{i}|u_{i-1})}$
Since $e^{\delta A|X_{i}|}$ is increasing in $X_{i}$ (as in K-SAT problem), and $m_{i}(u_{i-1}+X_{i}/N)Z_{i}(u_{i-1}+X_{i}/N)$ is decreasing in $X_{i}$ for fixed $u_{i}$, the negative association (3.1) implies,
$\mathbb{E}_{i-1}^{\!{}^{(M)}}[\,e^{\delta A|X_{i}|}\,]$ $\leq\frac{\displaystyle\int e^{\delta A|X_{i}|}\ P(dX_{i}|u_{i-1})\int m_{i}(u_{i-1}+X_{i})Z_{i}\ P(dX_{i}|u_{i-1})\;}{\displaystyle\int m_{i}(u_{i-1}+X_{i})Z_{i}\ P(dX_{i}|u_{i-1})}\leq K$
Thus, the exponential-moment condition in Lemma 3.1, inequality (3.6), holds. Now, similar to Theorem 2.2, we only need to prove
$\big|\,\mathbb{E}^{\,\!{}^{(M)}}(X_{\ell}\mid u_{i})-\mathbb{E}^{\,\!{}^{(M)}}(X_{\ell}\mid u_{i}^{\prime})\,\big|\,\leq\,L|u_{i}-u_{i}^{\prime}|~~~~for~i<l$
i.e. $E^{(M)}_{i}(X_{\ell})$ is Lipschitz continuous function of $u_{i}$.
Recall that in the proof of Theorem 2.2, smooth $\mu_{i}(u_{i})$ and $\nu_{i}(u_{i})$ are employed to prove “uniqueness of solution” (in RW “local” smoothness is “global” smoothness because it won’t change with generations—steps). In the case of BRW, we need smoothness of $\mu_{i}^{\,\!{}^{(M)}}(u_{i}):=E_{i}^{\,\!{}^{(M)}}(X_{i+1}|u_{i})$, $\nu_{i}^{\,\!{}_{(M)}}(u_{i}):=E_{i}^{\,\!{}^{(M)}}(X_{i+1}^{2}|u_{i})$, and of $\mathbb{E}_{i}^{(M)}\!\big(m(u_{i+M})\mid u_{i}\big)$. To this end, we first transform the “local” smoothness (given as hypotheses in the theorem) into “global” smoothness; i.e. to prove $\mu_{i}^{\,\!{}^{(M)}}(u_{i})$, $\nu_{i}^{\!{}^{(M)}}(u_{i})$ and $\beta_{i}^{\,\!{}^{(M)}}(u_{i})$ are smooth functions of $u_{i}$. We proceed by induction on the generations, recalling that the hypotheses in the theorem, the assumptions on $\beta_{i}$, $\mu_{i}$ and $\nu_{i}$ (without “(M)”), ensure that the base case (generation 0) holds. The following concerns generation $M$, where we use $n$ instead of $i$ for easy of comparison with the proof of Theorem 2.2.
Given birthplace $u_{n}$, its descendant population is
$\begin{split}&Z_{n}^{(M)}(u_{n})=Z_{n}^{(M)}(u_{n-1}+\frac{X_{n}}{N})\\ &=\int\Big(\prod_{i=1}^{M}m_{n+i}\Big)\ P(dX_{n+1}\cdots dX_{n+M}\mid u_{n}),\quad m_{i}=m(u_{i-1})\\ &=\int\Big(\prod_{i=1}^{M-1}m_{n+i}\Big)\ P(dX_{n+1}\cdots dX_{n-1+M}\mid u_{n})\int m_{n+M}P(dX_{n+M}\mid u_{n-1+M})\\ &=\int\Big(\prod_{i=1}^{M-1}m_{n+i}\Big)\ P(dX_{n+1}\cdots dX_{n-1+M}\mid u_{n})\,m_{n+M}\\ &=Z_{n}^{(M-1)}(u_{n})\cdot\frac{\displaystyle\int\Big(\prod_{i=1}^{M-1}m_{n+i}\Big)\ P(dX_{n+1}\cdots dX_{n-1+M}\mid u_{n})\,m(u_{n+M-1})}{\displaystyle Z_{n}^{(M-1)}(u_{n})}\\[5.0pt] &=Z_{n}^{(M-1)}(u_{n})\cdot\mathbb{E}_{n}^{\,\!{}^{(M-1)}}\left[\,m(u_{n+M-1})\,\right]\\[5.0pt] &=Z_{n}^{(M-2)}(u_{n})\cdot\mathbb{E}_{n}^{\,\!{}^{(M-2)}}\left[\,m(u_{n+M-2})\,\right]\mathbb{E}_{n}^{\,\!{}^{(M-1)}}\left[\,m(u_{n+M-1})\,\right]\\ &\vdots\\ &=\prod_{k={0}}^{M-1}\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right],\quad\,\mathbb{E}_{n}^{(0)}[\,m(u_{n})\,]=m(u_{n})\end{split}$ (3.8)
where
$\mathbb{E}_{n}^{{(k)}}\left[\,m(u_{n+k})\,\right]=\frac{\displaystyle\int m_{n+1}\cdots m_{{n+k}}\ P(dX_{n+1}\cdots dX_{n+k}|u_{n})\,m(u_{n+k})}{\displaystyle\int m_{n+1}\cdots m_{{n+k}}\ P(dX_{n+1}\cdots dX_{n+k}|u_{n})}$
Similarly,
$Z_{n+1}^{(M-1)}(u_{n+1})=\int\Big(\prod_{i=n+2}^{n+M}m_{i}\Big)\ P(dX_{n+2}\cdots dX_{n+M}\mid u_{n+1})=\prod_{k={0}}^{M-2}\mathbb{E}_{n+1}^{(k)}\left[\,m(u_{n+1+k})\,\right]$
and thus,
$Z_{n+1}^{(M-1)}(u_{n})=\prod_{k={0}}^{M-2}\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})\mid u_{n}\!+\!0)\,\big],\quad\,\mathbb{E}_{n+1}^{(0)}\big[\,m(u_{n}+0)\,\big]=m(u_{n})$ (3.9)
where
$\mathbb{E}_{n+1}^{{(k)}}\left[\,m(u_{n+1+k})\,\right]=\frac{\displaystyle\int m_{n+2}\cdots m_{{n+k}}\ P(dX_{n+2}\cdots dX_{n+k}|u_{n}+0)\,m(u_{n+k})}{\displaystyle\int m_{n+2}\cdots m_{{n+k}}\ P(dX_{n+2}\cdots dX_{n+k}|u_{n}+0)}$
Given birthplace $u_{n+1}$($u_{n}+{X_{n+1}}/{N})$, $\mathbb{E}_{n+1}^{(k)}\left[\,m(u_{n+1+k})\,\right]$ is a function of $u_{n+1}$. Let
$\mathbb{E}_{n+1}^{(k)}\left[\,m(u_{n+1+k})\,\right]:=f_{k}\left(u_{n}+\frac{X_{n+1}}{N}\right)$
Then,
$\mathbb{E}_{n+1}^{(k)}\left[\,m(u_{n+1+k})\mid X_{n+1}=0\,\right]=f_{k}\left(u_{n}\right)$
Define
$\begin{split}\mu^{(M)}_{n}(u_{n}):&=\mathbb{E}_{n}^{(M)}\big[\,X_{n+1}|u_{n}\,\big]\\ &=\frac{\displaystyle\int X_{n+1}\cdot m_{n+1}\cdots m_{n+M}\,P(dX_{n+1}\cdots dX_{n+M}|u_{n})}{\displaystyle\int m_{n+1}\cdots m_{n+M}\,P(dX_{n+1}\cdots dX_{n+M}|u_{n})},\quad m_{n+1}=m(u_{n})\\ &=\frac{\displaystyle\int X_{n+1}\cdot m(u_{n})\,P(dX_{n+1})\,\int\Big(\prod_{n+2}^{n+M}m_{i}\Big)\,P(dX_{n+2}\cdots dX_{n+M}|u_{n+1})}{Z_{n}^{(M)}(u_{n})}\\ &=\frac{\displaystyle\int X_{n+1}\cdot m(u_{n})\,P(dX_{n+1})\,Z_{n+1}^{(M-1)}(u_{n+1})}{Z_{n}^{(M)}(u_{n})}\\ &=\frac{\displaystyle\int X_{n+1}m(u_{n})\,P(dX_{n+1})\frac{Z_{n+1}^{(M-1)}(u_{n+1})}{Z_{n+1}^{(M-1)}(u_{n}+0)}}{Z_{n}^{(M)}(u_{n})/Z_{n+1}^{(M-1)}(u_{n}+0)}\end{split}$ (3.10)
(In RW, $m_{n+1}m_{n+2}\cdots=1$ and $\mu^{(M)}_{n}(u_{n})=\mu_{n}(u_{n})$ )
$\overline{X}_{n+1}^{(1)}:$ $=\frac{\displaystyle\int X_{{n+1}}m(u_{{n}})\,P(dX_{{n+1}}\mid u_{{n}})}{\displaystyle\int m(u_{{n}})\,P(dX_{{n+1}}\mid u_{{n}})}=\int X_{{n+1}}\,P(dX_{{n+1}}\mid u_{{n}})$
In order to estimate $Z_{n}(u_{n})/Z_{n+1}(u_{n})$, we investigate ${\mathbb{E}_{n}^{(k+1)}[\,\cdot\,]}/{\mathbb{E}_{n+1}^{(k)}[\,\cdot\,]}$. Let $Z_{n+1}$ denote
$Z_{n+1}:=\int\Big(\prod_{i=n+2}^{n+k}m_{i}\Big)\,P(dX_{n+2}\cdots dX_{n+k}\mid u_{n+1}),\quad m_{i}=m(u_{i-1}).$
Then, the average birth rate of the the latest generation, $k$ generations after the $n\text{th}$, is
$\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]=\frac{\displaystyle\int m_{n+1}m_{n+2}\cdots m_{{n+k}}\,P(dX_{n+1}\cdots dX_{{n+k}}|u_{n})\,m(u_{n+k})}{\displaystyle\int m_{n+1}m_{n+2}\cdots m_{{n+k}}\,P(dX_{n+1}\cdots dX_{n+k}|u_{n})\quad\big(\!:=Z_{n}(u_{n})\big)}$
$=\frac{\displaystyle\int m_{n+1}P(dX_{n+1}|u_{n})\int\Big(\prod_{i=n+2}^{n+k}m_{i}\Big)\,P(dX_{n+2}\cdots dX_{n+k}|u_{n+1})\,m(u_{n+k})}{\displaystyle Z_{n}(u_{n})}$
$=\frac{\displaystyle\int m_{n+1}P(dX_{n+1}|u_{n})Z_{n+1}\frac{\int\Big(\prod_{i=n+2}^{n+k}m_{i}\Big)\,P(dX_{n+2}\cdots dX_{n+k}|u_{n+1})\cdot m(u_{n+k})}{\displaystyle Z_{n+1}}}{\displaystyle Z_{n}(u_{n})}$
$=\frac{\displaystyle\int m_{n+1}P(dX_{n+1}|u_{n})Z_{n+1}\,\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right]}{Z_{n}(u_{n})}$
$=\frac{\displaystyle\int m_{n+1}P(dX_{n+1}|u_{n})\left[\int\Big(\prod_{i=n+2}^{n+k}m_{i}\Big)\,P(dX_{n+2}\cdots dX_{n+k}|u_{n+1})\right]\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right]}{Z_{n}(u_{n})}$
$=\frac{\displaystyle\int\Big(\prod_{i=n+1}^{n+k}m_{i}\Big)\,P(dX_{n+1}...dX_{n+k}|u_{n})\,\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right]}{\displaystyle\int\Big(\prod_{i=n+1}^{n+k}m_{i}\Big)\,P(dX_{n+1}...dX_{n+k}\mid u_{n})}$
$=\mathbb{E}_{n}^{(k)}\left[\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right]\right]$
That is, for all $n$ and $k$
$\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]=\mathbb{E}_{n}^{(k)}\left[\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right]\right]\quad$ (3.11)
The remainder of the proof is conduced in the following order. After the base case ($k=0$), we show that for $k=1$, $\overline{X}_{n+1}^{(1)}$ is continuous, where
$\overline{X}_{n+1}^{(1)}:=\mathbb{E}_{n}^{{(1)}}(X_{n+1})=\frac{\displaystyle\int X_{n+1}m_{n+1}P(dX_{n+1}|u_{n})}{\displaystyle\int m_{n+1}P(dX_{n+1}|u_{n})}=\int X_{n+1}P(dX_{n+1}|u_{n})$
We next prove $\mathbb{E}_{n}^{(1)}\left[\,m(u_{n+1})\,\right]$ is continuous in $u_{n}$, using the continuity of $\overline{X}_{n+1}^{(1)}$. We then show continuity of $\overline{X}_{n+1}^{(2)}$ using $\mathbb{E}_{n}^{(1)}\left[\,m(u_{n+1})\,\right]$, and so on:
$\overline{X}_{n+1}^{(1)}\;\Rightarrow\;\mathbb{E}_{n}^{(1)}\;\Rightarrow\;\overline{X}_{n+1}^{(2)}\;\Rightarrow\;\mathbb{E}_{n}^{(2)}\;\Rightarrow\;\cdots$
leading to continuity of $\mu^{(M)}_{n}(u_{n})$ and $\mathbb{E}_{n}^{(M)}\left[\,m(u_{n+M})\,\right]$ for all $M\lesssim N$. The continuity of $\nu_{n}^{{(k)}}(u_{n})$ can be proved given $\mu^{(k)}_{n}(u_{n})$ and $\mathbb{E}_{n}^{(k-1)}\left[\,m(u_{n+k-1})\,\right]$.
When $k=1$, for any $i$
$\mathbb{E}_{i}^{(1)}\left[\,m(u_{i}+X/N)\,\right]$ $=\frac{\displaystyle\int m(u_{i})\,P(dX|u_{i})\,m\Big(u_{i}+\frac{X}{N}\Big)}{\displaystyle\int m(u_{i})\,P(dX|u_{i})}=\int m\Big(u_{i}+\frac{X}{N}\Big)P(dX|u_{i})$
is a smooth function of $u_{i}$ (meaning it has continuous derivatives up to some required order) since $m(x)$ is uniformly smooth in $x$ and the conditional distribution $P(dX_{i+1}|u_{i})$ varies smoothly with $u_{i}$. For instance,
$\mathbb{E}_{n}^{(1)}\left[\,m\Big(u_{n}+\frac{X_{n+1}}{N}\Big)\,\right]=m(u_{n})+\frac{m^{\prime}(u_{n})}{N}\overline{X}_{n+1}^{(1)}+O(N^{-2})$ (3.12)
Let $f_{i}^{(0)}(u_{i})$ denote $m(u_{i})$ and write
$\overline{X}_{n+1}^{(k)}=\mathbb{E}_{n}^{(k)}(X_{n+1}),\quad f_{n+1}^{(k-1)}(u_{n+1})=\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right];$
thus (3.12) can be written as
$f_{n+1}^{(1)}(u_{n+1})=f_{n}^{(0)}(u_{n})+\frac{{f_{n}^{(0)}}^{\prime}(u_{n})}{N}\overline{X}_{n+1}^{(1)}+O(N^{-2})$ (3.13)
and (3.11) becomes
$f_{n}^{(k)}(u_{n})=\mathbb{E}_{n}^{(k)}\left[f_{n+1}^{(k-1)}(u_{n+1})\right]$
Now assume for $k\geq 2$, $f^{(k-1)}(u_{n+1})$ is smooth in $u_{n+1}$ for any $n+1$. Employing a Taylor expansion, we have
$\begin{split}f_{n}^{(k)}(u_{n})&=\mathbb{E}_{n}^{(k)}\!\left[f_{n+1}^{(k-1)}\Big(u_{n}+\frac{X_{n+1}}{N}\Big)\right]\\ &=f_{n+1}^{(k-1)}\!\left(u_{n}\right)+\frac{\overline{X}_{n+1}^{(k)}}{N}{f_{n+1}^{(k-1)}}^{\prime}(u_{n})+O\!\left(N^{-2}\right)\end{split}$ (3.14)
which is smooth in $u_{n}$. Observe that the term $\frac{\overline{X}_{n+1}^{(k)}}{N}{f_{n+1}^{(k-1)}}^{\prime}(u_{n})$ involves no second order derivatives of $f_{n+2}^{(k-2)}$ which is absorbed into the $O(N^{-2})$ term. Thus, we have proved that for any $k$, $\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]$ is smooth in $u_{n}$.
$\overline{X}_{n+1}^{(k)}:=\mathbb{E}_{n}^{(k)}[\,X_{n+1}\,\mid u_{n}]=\frac{\displaystyle\int X_{n+1}\,m_{n+1}m_{n+2}\cdots m_{k}\,P(dX_{n+1}...dX_{k}\mid u_{n})}{\displaystyle\int m_{n+1}m_{n+2}\cdots m_{k}\,P(dX_{n+1}\cdots dX_{k}\mid u_{n})}$ (3.15)
$\frac{\mathbb{E}_{n}^{(k+1)}\left[\,m(u_{n+1+k})\,\right]}{\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})|X_{n+1}=0\,\big]}$ $=\frac{\displaystyle f_{n+1}^{(k)}\!\left(u_{n}\right)+\frac{\overline{X}_{n+1}^{(k+1)}(u_{n})}{N}{f_{n+1}^{(k)}}^{\prime}\!\left(u_{n}\right)+O\!\left(N^{-2}\right)}{f_{n+1}^{(k)}\left(u_{n}+0\right)}$
$=1+\frac{\overline{X}_{n+1}^{(k+1)}(u_{n})}{N}\cdot\frac{{f_{n+1}^{(k)}}^{\prime}\!\left(u_{n}\right)}{f_{n+1}^{(k)}\left(u_{n}\right)}+O\!\left(N^{-2}\right)$
Notice that $Z_{n+1}(u_{n})$ is $Z_{n+1}\big(u_{n}+{X_{n+1}}/{N}\big)|_{X_{n+1}=0}$ . From (3.8) and (3.9), we have
${Z_{n}^{(M)}(u_{n})}/{Z_{n+1}^{(M-1)}(u_{n})}$ $=\frac{\displaystyle\prod_{k=0}^{M-1}\mathbb{E}_{n}^{(k)}\big[\,m(u_{n+k})\,\big]}{\displaystyle\prod_{k=0}^{M-2}\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})|X_{n+1}=0\,\big]}$
$=\frac{\displaystyle m(u_{n})\prod_{k=0}^{M-2}\;\mathbb{E}_{n}^{(k+1)}\big[\,m(u_{n+1+k})\,\big]}{\displaystyle\prod_{k=0}^{M-2}\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})|X_{n+1}=0\,\big]}$
$=m(u_{n})\,\prod_{k=0}^{M-2}\left(1+\frac{\overline{X}_{n+1}^{(k+1)}(u_{n})}{N}\cdot\frac{{f_{n+1}^{(k)}}^{\prime}\!\left(u_{n}\right)}{f_{n+1}^{(k)}\left(u_{n}\right)}+O\!\left(N^{-2}\right)\right)$
where $\overline{X}_{n+1}^{(k+1)}:=\mathbb{E}_{n}^{(k+1)}\left[\,X_{n+1}\,\mid u_{n}\right]$, and
$Z_{n+1}(u_{n+1})/Z_{n+1}(u_{n})$ $=\frac{\displaystyle\prod_{k=0}^{M-2}\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})\,\big]}{\displaystyle\prod_{k=0}^{M-2}\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})|X_{n+1}=0\,\big]}$
$=\prod_{k=0}^{M-2}\frac{\displaystyle f_{n+1}^{(k)}\!\left(u_{n}\right)+{f_{n+1}^{(k)}}^{\prime}\!\left(u_{n}\right)\frac{X_{n+1}}{N}+O\!\left(N^{-2}\right)}{\displaystyle f_{n+1}^{(k)}\left(u_{n}\right)}$
$=\prod_{k=0}^{M-2}\left(1+\frac{X_{n+1}}{N}\frac{{f_{n+1}^{(k)}}^{\prime}(u_{n})}{f_{n+1}^{(k)}\left(u_{n}\right)}+O\!\left(N^{-2}\right)\right)$
Then, (3.10)—that is, $\mu^{(M)}_{n}(u_{n})$—becomes, neglecting terms of order $O\!\left(N^{-2}\right)$,
$\overline{X}_{n+1}^{(M)}(u_{n})$ $=\int\!X_{n+1}\prod_{k=0}^{M-2}\!\left(1+\frac{1}{N}\frac{{f_{n+1}^{(k)}}^{\prime}(u_{n})}{f_{n+1}^{(k)}(u_{n})}\Big(X_{n+1}-\overline{X}_{n+1}^{(k+1)}(u_{n})\Big)\right)\!P(dX_{n+1}|u_{n})$
This is a smooth function of $u_{n}$ since $f_{n+1}^{(k)}$, $f_{n+1}^{{(k)}}$$^{\prime}$ and $\overline{X}_{n+1}^{(k+1)}(u_{n})$ are smooth by induction and $f_{n+1}^{(k)}$ are bounded (recall $m(u_{i})$ is bounded). Similarly it can be shown $Z_{n+1}(u_{n+1})/Z_{n+1}(u_{n})$ is a smooth function of $u_{n}$, conditinal on $X_{n+1}$. Therefore $\mu^{(M)}_{n}(u_{n})$ is a smooth function of $u_{n}$; see (3.10) for $\mu^{(M)}_{n}(u_{n})$.
In the same way, we can prove the smoothness of $\nu_{n}^{{(M)}}(u_{n})$ and $\mathbb{E}_{n}^{(M)}\left[\,m(u_{n+M})\,\right]$.
The rest of the proof is the same as in Theorem 2.2. It is worthy noting that to complete the proof we need to use
[…]
This in turn requires concentration inequality for $x_{1}+x_{2}+...x_{j}$ which holds by induction. ∎
Thus we have shown that the negative association assumption (3.1) is sufficient to establish concentration inequalities for BRW, if the RW is not too scattering. Is it a necessary condition? Now consider the classical random walk of coin tossing, $S_{n}=X_{1}+...X_{n}$ with $X_{i}\in{1,-1}$ (so $E(S_{n})$ = 0), which is also a BRW with branching factor $m_{i}=1$. It is easy to check that if the branching factor is $m=1+\delta|u_{i}|$ for $\delta>0$, then the BRW is scattering. Therefore in a degree the negative association assumption is also a necessary condition.
In the situation of bounded $X_{i}$, a particle always stays around the mean trajectory as long as its birthplace $u_{i-1}$ is near that trajectory. In the case of unboundedness, however, a particle may move into regions near singularities (e.g., critical points in K–SAT) where $u_{i}$ is not smooth. In such regions, $\mathbb{E}(e^{\delta Xi})<K$ fails, among others. The good news is, because concentration inequality holds in the neighborhood area of the mean trajectory—is at least $\epsilon$ far away from the singularity—the outliers account for a negligible proportion of the population and can be ignored. Crucially, pruning these rare extremal paths does not weaken the concentration. The following theorem formalizes this observation.
Theorem 3.2.
Let $\epsilon$ be small compared with $\delta$, $N^{1/2}\ll\lambda<\epsilon N$, and as before,
$S_{i}:=\sum_{t=1}^{i}X_{t},\quad\Delta S_{i}:=|S_{i}-\mathbb{E}S_{i}|,\quad u_{i}:=\frac{S_{i}}{N}$
Define the “good neighborhood” around the mean path
$G_{i-1}:=\Big(\bar{u}_{i-1}-\frac{\lambda}{N},\bar{u}_{i-1}+\frac{\lambda}{N}\Big),\quad\bar{u}_{i-1}:=\frac{\mathbb{E}S_{i-1}}{N}$
Assume as before
$E\left[e^{\delta|X_{i}|}\,\big|\,u_{i-1}\in G_{i-1}\right]\leq K\quad\text{{(bounded MGF)}}$
Then
$\Pr\Big(\max_{1\leq i\leq n}\Big|\,S_{i}-\mathbb{E}[S_{i}]\,\Big|\leq\lambda\Big)\geq\prod_{i=1}^{n}\Big(1-2e^{-{c\lambda^{2}}/{i}}\Big)$ (3.16)
In our context, $N^{1/2}\ll\lambda<\epsilon N$ and $n<N$. Thus
$\Pr\Big(\displaystyle\max_{1\leq i\leq n}\Big|\,S_{i}-\mathbb{E}[S_{i}]\,\Big|\geq\lambda\Big)\leq 2n\exp\left(-\frac{c\lambda^{2}}{n}\right)$
which is extremely (exponentially) small.
Proof of (3.16).
The proof proceeds by induction. For each $i\geq 1$, set
$A_{i}:={\Delta S_{i}\leq\lambda}$
Then, the following equivalence holds:
$\max_{1\leq i\leq n}\big|\,S_{i}-\mathbb{E}[S_{i}]\,\big|\leq\lambda\Longleftrightarrow A_{1}\wedge A_{2}\wedge\cdots\wedge A_{n}$
Note the implications
$A_{i-1}$ $\iff|S_{i-1}-\mathbb{E}S_{i-1}|\leq\lambda$
$\iff\left|u_{i-1}-\bar{u}_{i-1}\right|\leq\frac{\lambda}{N}$
$\iff u_{i-1}\in G_{i-1}$
For the base case, we have $u_{0}\in G_{0}$ where
$G_{0}:=\Big(u_{0}-\frac{\lambda}{N},u_{0}+\frac{\lambda}{N}\Big);$
bounded MGF is satisfied for $i=1$. Thus concentration inequality holds,
$\Pr\Big(|\,S_{1}-\mathbb{E}[S_{1}]\,|\leq\lambda\Big)\geq\Big(1-2e^{-{c\lambda^{2}}/1}\Big)$
Assume the claim (3.16) holds for all $i$ through to n1n-1, i.e,
$\Pr\Big(A_{1}\wedge A_{2}\wedge\cdots\wedge A_{n-1}\Big)\geq\prod_{i=1}^{n-1}\Big(1-2e^{-{c\lambda^{2}}/{i}}\Big)$
Then, given
$u_{1}\in G_{1},u_{2}\in G_{2},...,u_{n-1}\in G_{n-1}$
the standard concentration inequality can be established for $S_{n}$ by the same argument as in Lemma 2.3 and Theorem 2.2 :
$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\geq\lambda\,\Big|\,u_{1}\in G_{1},...,u_{n-1}\in G_{n-1}\Big)\leq 2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
so
$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\leq\lambda\,\Big|\,u_{1}\in G_{1},...,u_{n-1}\in G_{n-1}\Big)\geq 1-2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
Thus
$\Pr\Big(A_{1}\wedge\cdots\wedge A_{n-1}\wedge A_{n}\Big)$ $=\Pr\Big(A_{n}\,\big|\,A_{1}\wedge\cdots\wedge A_{n-1}\Big)\cdot\Pr\Big(A_{1}\wedge\cdots\wedge A_{n-1}\Big)$
$\geq\Big(1-2e^{-c\frac{\lambda^{2}}{n}}\Big)\prod_{i=1}^{n-1}\Big(1-2e^{-{c\lambda^{2}}/{i}}\Big)$
The claim follows. ∎
4 Applications
See [31]

### v5

@@
 While classical concentration inequalities are typically restricted to two special cases—independence and martingale difference sequences—we extend concentration inequalities to a much broader class of stochastic processes by relaxing these foundational conditions.
-Specifically, heuristically and in the language of calculus, while independence and the martingale difference property correspond to
+Specifically, in calculus terms, while independence and the martingale difference property correspond respectively to
 $\frac{\partial y}{\partial t}=\text{constant},\quad\displaystyle\frac{\partial y}{\partial t}=0$
-respectively, we relax these conditions to
+we relax these conditions to
 a general setting that requires only the existence of a drift $\frac{\partial y}{\partial t}$ which is allowed to vary with the past state.
-Furthermore, concentration inequalities are established for branching random walks.
+Furthermore, concentration inequalities for branching random walks are established.
 Branching random walk,
@@
 $S_{n}$ can be formulated as $S_{n}=X_{1}+X_{2}+...+X_{n}$; $X_{i}=S_{i}-S_{i-1}$.
-If some conditions are satisfied, among which independence is crucial, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT).
-For $S_{n}$ beyond the neighborhood, i.e., $|S_{n}-E(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [23]), also referred as Chernoff bound,
-$Pr(|S_{n}-E(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
+Under certain conditions, in particular independence, $S_{n}$ converges to a Gaussian distribution in a $\sqrt{n}$ neighborhood of its mean — central limit theorem (CLT).
+For $S_{n}$ beyond the neighborhood, i.e., $|S_{n}-\mathbb{E}(S_{n})|\gg\sqrt{n}$, Chernoff inequality ([14] [24]), also referred as Chernoff bound,
+$\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda)\leq e^{-c\frac{\lambda^{2}}{n}}$ (1.1)
 tells us how unlikely this occurrence is.
@@
 [19], [4], [40], [31]).
-However, for a long time, there had been no significant improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [40].
+For a long time, there had been no substantive improvement in the eliminating boundedness condition (the trivial idea of truncation gives weak results, though) [^2] In fact, it was not until 1995 that the first improvement in Hoeffding inequality was made [42].
 The reason might be two folds.
 First, there lacked motivation.
-Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results concealed the essential reason why the inequalities hold.
+Second, nifty techniques used in the pioneer work and the later literatures as well to achieve finer results somehow concealed the essential reason why the inequalities hold..
 In early 2000s [39] relaxed the boundedness to finite exponential moments; i.e.
-for a $\delta>0,E(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-E(S_{n})|\geq\lambda|)\leq e^{-\sim\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$ — not Chernoff bound.
-Later, motivated by the study of the free energy of a directed polymer in a random environment, [32] gave Chernoff bound under the hypothesis that, for a $\delta>0,E(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$).
-Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [25].
-No attempt is known to get rid of martingale condition.
-The first major contribution of this paper is in this direction.
-1.2 Notation and contribution
+for a $\delta>0,\mathbb{E}(e^{\delta X_{i}})$ is uniformly bounded for all $i$, and gave $\Pr(|S_{n}-\mathbb{E}(S_{n})|\geq\lambda|)\leq e^{-c\big(\frac{\lambda^{2}}{n}\big)^{1/3}}$— not Chernoff bound.
+Later, motivated by the study of the free energy of a directed polymer in a random environment, [34] gave Chernoff bound under the hypothesis that, for a $\delta>0,\mathbb{E}(e^{\delta X_{i}}\mid\mathcal{F}_{i-1})$ is bounded for all $i$ (where $\mathcal{F}_{i-1}$ is a filtration for the martingale difference sequence $X_{i}$).
+Other improvements of concentration inequalities in relaxing boundedness condition can be found in [8] [20] [26] [21].
+However, to our knowledge, no existing work eliminates both the martingale and boundedness conditions while achieving sub-Gaussian concentration; related work by Wormald (1995) is discussed later.
+The first major contribution of this paper is to do so.
+Informally, the small per-step effect in a random process can be controlled by the Lipschitz continuity of the mean increment with respect to the history, not necessarily by “almost independence”.
+1.2 Notation and contributions
 Let $U_{i}$ denote the partial sum of the first $i$ terms of $S_{n}=\sum_{j=1}^{n}X_{j}$, namely, $U_{i}=X_{1}+X_{2}+...+X_{i}$.
@@
 For the sake of readability, in some places we use $Y_{n}$ for $S_{n}$.
-We let $\mathbb{E}_{i}[\,\cdot\,]$ to abbreviate $\mathbb{E}[\,\cdot\mid u_{i}\,]$.
+We let $\mathbb{E}_{i}[\,\cdot\,]$ abbreviate $\mathbb{E}[\,\cdot\mid u_{i}\,]$.
 Note that $\mathbb{E}_{i}[X_{t}]$ is a function of both $t$ and past state $(u_{i})_{i\leq t}$.
@@
 where $S_{n}=X_{1}+X_{2}+...+X_{n}$, $p_{1}(x_{1})p_{2}(x_{2})...p_{n}(x_{n})$ is the probability density of $(x_{1},...x_{n})$ (here and throughout we use $p_{i}(x_{i})$ for $p_{i}(x_{i}|u_{i-1})$ for notational simplicity).
-In algorithm analysis and random graphs study, which are described as graph evolution stochastic process (see e.g.
+In algorithm analysis and random graphs study, which are modelled as graph evolution stochastic processes (see e.g.
 [36], [44], [45], [46], [47] and references therein), the differential equation method (DEM) started in 80’s.
 Recently DEM has obtained some important results and received increasingly attention.
-The procedure of DEM consists of (a) verifying that the random parameters concerned concentrate around their means, and (b) finding solutions to the differential equations satisfied the means.
+The procedure of DEM (see [44] for a general framework, and [28] for its stronger version) consists of (a) verifying that the random parameters concerned concentrate around their means, and (b) finding solutions to the differential equations satisfied the means.
 Concentration inequalities (e.g.
@@
 For the applications, brief list of the results obtained is presented in this paper; for details, the readers are referred to [33], a companion paper.
-2 An extension of Heoffding-Azuma inequality
-For warming up, let’s present a “downgraded” proof for Azuma’s inequality.
-Possibly this proof could be found in the Azuma’s trash can.
-Let’s first introduce a simple and elegant inequality due to [32] and [41].
-Given here is the version from [41] which improves [32].
+2 Extension of Hoeffding-Azuma inequality
+As a pedagogical warm-up, we present a “downgraded” proof for Azuma’s inequality. To start with, we introduce a simple inequality following [34] and [43].
 Lemma 2.1.
@@
 (Azuma’s Inequality) Let $\{X_{i}\}_{i=1}^{n}$ be a sequence of martingale difference, with $|X_{i}|\leq 1$ (without loss of generality), and $S_{n}=X_{1}+X_{2}+...+X_{n}$, then
-$Pr(|S_{n}-ES_{n}|\geq\lambda)\leq 2\exp\Big(-\frac{\lambda^{2}}{2nc^{2}}\Big)$
+$\Pr(|S_{n}-\mathbb{E}S_{n}|\geq\lambda)\leq 2\exp\Big(-\frac{\lambda^{2}}{2nc^{2}}\Big)$
 To prove the inequality, Azuma ([2]) employed the following inequalities, for $|X|\leq c$,
@@
 $\exp({\sim nt^{2}-t\lambda})$
-which is exponentially small under the choice $t\sim n^{-1/2}$, even for $\lambda\ll n$.
-Proof (Downgraded version of Azuma’s inequality).
-We only consider $\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$, as the proof of the reverse $\mathbb{P}(S_{n}-\mathbb{E}S_{n}\leq-\lambda)$ is exactly the same.
+which, given $\lambda=o(n)$, is exponentially small, $\exp(-n^{p})$ with $0<p<1$, under the choice $t\sim n^{(p-1)/2}$.
+Proof (A non-sharp version of Azuma’s inequality).
+We only consider ${\Pr}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$, as the proof of the reverse ${\Pr}(S_{n}-\mathbb{E}S_{n}\leq-\lambda)$ is exactly the same.
 By Markov’s inequality, inductive application of Lemma 2.1 and optimization of $t$, we have
-$\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$ $\leq e^{-t\lambda}\cdot E\{e^{tS_{n-1}}\cdot\mathbb{E}(e^{tX_{n}}\mid X_{1},...,X_{n-1})\}$
+${\Pr}(S_{n}-\mathbb{E}S_{n}\geq\lambda)$ $\leq e^{-t\lambda}\cdot E\Big[e^{tS_{n-1}}\cdot\mathbb{E}(e^{tX_{n}}\mid X_{1},...,X_{n-1})\Big]$
 […]
-$\leq e^{-t\lambda}\cdot\mathbb{E}\{e^{tS_{n-1}}\cdot e^{t^{2}e^{c}}\}~~~~~~~\mbox{(}|X_{i}|\leq c\mbox{)}$
+$\leq e^{-t\lambda}\cdot\mathbb{E}\Big[e^{tS_{n-1}}\cdot e^{t^{2}e^{c}}\Big]~~~~~~~\mbox{(}|X_{i}|\leq c\mbox{)}$
 $\leq\cdots\leq\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}\quad\mbox{(cf. Azuma's $\exp{\left(\displaystyle-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$)}$
-The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [32] and [41]) without compromising strength of results; from the practical point of view, the difference between $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ and $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$ is immaterial – both are square-exponential decay.
-In addition, it is demonstrated that nifty techniques are not necessary for obtaining concentration inequalities.
-Next, we will show what else can make the concentration inequality hold, other than independence and martingale difference. To this end, define Doob’s (or McDiarmid’s) martingale
+The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [34] and [43]) without compromising strength of results; from the practical point of view, the difference between the constant factors in the exponent, i.e., $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ versus $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$, is immaterial – both are square-exponential decay.
+In addition, it is shown that sub-Gaussian tail bounds can be obtained without “nifty” techniques.
+Next, we show what weaker conditions can ensure concentration inequalities, beyond independence, martingale differences, and boundedness. To this end, define Doob’s (or McDiarmid’s) martingale
 $d_{i}$ $=\mathbb{E}_{i}(X_{i}+\cdots+X_{n})-\mathbb{E}_{i-1}(X_{i}+\cdots+X_{n})$ (2.4)
 It is easy to check
-$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\quad$
-$E_{i-1}(d_{i})=0$. Notice
+$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\,,\quad E_{i-1}(d_{i})=0$
+Note
 $d_{i}$ $=X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)$ (2.5)
@@
 $|\mathbb{E}_{i}(X_{\ell})-\mathbb{E}_{i-1}(X_{\ell})|\leq L\frac{|X_{i}|}{N}+L\frac{\mathbb{E}_{i-1}(\,|X_{i}|\,)}{N}$ (2.7)
-Note, $\mathbb{E}_{i}(X_{\ell})=\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
+Note, $\mathbb{E}_{i}(X_{\ell})\equiv\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
 Proof (Lemma 2.2).
@@
 $\mathbb{E}_{X_{i}^{\prime}}(\,\cdot\,)\equiv\mathbb{E}_{X_{i}^{\prime}}(\,\cdot\mid u_{i-1}),\quad X_{i}^{\prime}\sim\mathcal{L}(X_{i}\mid u_{i-1}).$
-With this notation,
+With this notation, we have
 $\bigl|\mathbb{E}_{i}\left(X_{\ell}\right)-\mathbb{E}_{i-1}\left(X_{\ell}\right)\bigr|$ $=\Bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{X_{i}^{\prime}}\!\left[\,\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\right]\Bigl|$
 $=\Bigl|\,\mathbb{E}_{X_{i}^{\prime}}\big[\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\big]\,\Bigl|$
-$\leq\mathbb{E}_{X_{i}^{\prime}}\Big(\,\bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\bigl|\,\Big)$
+$\leq\mathbb{E}_{X_{i}^{\prime}}\Big[\,\bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\bigl|\,\Big]$
 $\leq\mathbb{E}_{X_{i}^{\prime}}\!\left[\,L\frac{|X_{i}-X_{i}^{\prime}|}{N}\,\right],\quad\text{$\mathbb{E}_{i-1}\left[\,X_{\ell}\mid X_{i}\,\right]$ is L-Lipschitz in $X_{i}$}$
@@
 $\left|\,\mathbb{E}_{i-1}(X_{\ell}\mid X_{i})-\mathbb{E}_{i-1}(X_{\ell}\mid X_{i}^{\prime})\,\right|\leq L\displaystyle\frac{|X_{i}-X_{i}^{\prime}|}{N},\quad\text{for $\ell>i$}\;$ (2.8)
-and $\mathbb{E}[\,e^{\delta A|X_{i}|}\mid u_{i-1}\,]\leq K$ (exponential moment existence) for a constant $\delta>0$ , $i=1,2,$..., then for $n\leq M$
-$\Pr\big(\,|\,S_{n}-\mathbb{E}S_{n}\,|\geq\lambda\,\big)\leq 2\exp{\Big(\!-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\Big)}$ (2.9)
-Obviously, Poisson distribution $\text{Poisson}(\lambda_{i})$ meets the exponential moment condition.
+and $\mathbb{E}[\,e^{\delta A|X_{i}|}\mid u_{i-1}\,]\leq K$ (exponential moment existence) for a constant $\delta>0$ , $i=1,2,$..., then for $n\leq M$ $\dagger$$\dagger\daggerActually$, the lemma holds for $n=O(M)$; we assume $n\leq M$ here to simplify the argument.
+$\Pr\big(\,|\,S_{n}-\mathbb{E}S_{n}\,|\geq\lambda\,\big)\leq 2\exp{\Big(\!-c\frac{\lambda^{2}}{n}\Big)},\quad c=\frac{\delta^{2}}{4K^{2}}\;$ (2.9)
+Remarks.
+  • Condition (2.8) in Lemma 2.3 is equivalent to the following Lipschitz continuity
+    $\left|\mathbb{E}(X_{\ell}\mid u_{i})-\mathbb{E}(X_{\ell}\mid u_{i}^{\prime})\right|\leq L\left|u_{i}-u_{i}^{\prime}\right|\quad\mbox{for $i<l$}$ (2.10)
+    where
+    $u_{i}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}),\quad u_{i}^{\prime}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}^{\prime}).$
+  • For $n=O(N)$, the L-Lipschitz condtion may imply
+    $\left|\sum_{\ell=i+1}^{n}\mathbb{E}_{i-1}(X_{\ell}\mid X_{i})-\sum_{\ell=i+1}^{n}\mathbb{E}_{i-1}(X_{\ell}\mid X_{i}^{\prime})\right|\leq L\displaystyle{|X_{i}-X_{i}^{\prime}|}$
+    When $X_{i}$ and $X_{i}^{\prime}$ are bounded, this reduces to McDiarmid’s bounded difference condition [30] when applied to random walks.
+  • Independence and martingale difference assumptions both imply this L-Lipschitz condition (2.8). That is, the Lipschitz condition is weaker than those assumptions.
+  • Poisson, Binomial, and Geometric distributions, for example, all satisfy the finite exponential moment condition, as does the Gaussian distribution.
 Proof of Lemma 2.3.
 By definition (2.5),
-$\mathbb{E}_{i-1}\exp\left(\delta|d_{i}|\right)=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
+$\mathbb{E}_{i-1}\left[e^{\delta|d_{i}|}\right]=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
 $\leq\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|\mathbb{E}_{i-1}(X_{i})|+\delta\sum_{\ell=i+1}^{n}\big|\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big|\Big)\Big]$
@@
 Applying the exponential Markov inequality $\Pr[X\geq\lambda]\leq e^{-t\lambda}E(e^{tX})$ and noticing
-$E_{i-1}(d_{i})=0$
-we have
+$\mathbb{E}_{i-1}(d_{i})=0$
+it follows
 $\Pr\big(S_{n}-\mathbb{E}S_{n}\geq\lambda\,\big)$
@@
 Optimizing in $t$, we obtain
-$\mathbb{P}(S_{n}-\mathbb{E}S_{n}\geq\lambda)\leq\exp\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right).$
-For the other half side of the concentration inequality, we refer the fact that Lemma 2.1 holds for negative $t$ as well. The proof is exactly the same as above, which we omit. ∎
+${\Pr}(S_{n}-\mathbb{E}S_{n}\geq\lambda)\leq\exp\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)$
+For the other side of the concentration inequality, we refer to the fact that Lemma 2.1 holds for negative $t$ as well. The proof is identical to the above. ∎
 Remarks.
-  • Condition (2.8) in Lemma 2.3 is equivalent to the following Lipschitz continuity
-    $\left|\mathbb{E}(X_{\ell}\mid u_{i})-\mathbb{E}(X_{\ell}\mid u_{i}^{\prime})\right|\leq L\left|u_{i}-u_{i}^{\prime}\right|\quad\mbox{for $i<l$}$ (2.10)
-    where
-    $u_{i}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}),\quad u_{i}^{\prime}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}^{\prime}).$
-  • The condition of Lipschitz continuity (2.10) may be rephrased “uniqueness of solution”; i.e. small difference in initial position can not yield big difference in the trajectory.
+  • The condition of Lipschitz continuity (2.10) may be interpreted as “continuity” of the trajectory with respect to the history; i.e.
+small differences in initial position do not lead to vehement differences in the trajectory.
   • In the case of independence, $\mathbb{E}_{i}[X_{\ell}]=\mathbb{E}[X_{\ell}]$, and hence $\mathbb{E}_{i}[X_{\ell}]-\mathbb{E}_{i-1}[X_{\ell}]$ = 0.
@@
 In these cases, the condition of bounded difference or the existence of exponential moment alone suffices to imply the concentration inequalities.
-  • The condition of Lipschitz continuity (2.10) may be rephrased “uniqueness of solution”; i.e. small difference in initial position can not yield big difference in the trajectory.
   • Recall (1.3), the notations defined in the section of Introduction. In the notation of calculus, the condition (2.10) is
@@
 Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on.
-We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the region in question is distant enough from singularities).
+We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the domain in question lies sufficiently away from singularities).
 Theorem 2.2.
-If $E\big(e^{\delta A|X_{i}|}\mid u_{i-1}\big)\leq K$ (existence of an exponential moment), then in the region where $\mu_{i}$ and $\nu_{i}$ are smooth, the concentration inequality holds.
+If $\mathbb{E}\big(e^{\delta A|X_{i}|}\mid u_{i-1}\big)\leq K$ (existence of an exponential moment), then in the domain where $\mu_{i}$ and $\nu_{i}$ are smooth, the concentration inequality holds.
 In the notation of partial derivatives, the condition in Theorem 2.2 states that
-$\frac{\partial\mu(t,u)}{\partial u},\quad\frac{\partial^{2}\mu(t,u)}{\partial u^{2}},\quad\frac{\partial\nu(t,u)}{\partial u},\cdots$
-are uniformly bounded in the region of smoothness (i.e., away from singularities), in addition to the exponential moment condition.
+$\frac{\partial\mu(t,u)}{\partial u},\quad\frac{\partial^{2}\mu(t,u)}{\partial u^{2}},\quad\frac{\partial\nu(t,u)}{\partial u},\quad\cdots$
+are uniformly bounded in the domain of smoothness (i.e., away from singularities), in addition to the exponential moment condition.
 These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE).
@@
 Theorem 2.2 tells us, in terms of PDE of $\bar{u}$, what local behavior of $u$ can meet (2.10), for the concentration inequality to hold.
-Proof.
-We just need to prove Lipschitz continuity (2.10).
+Proof of Theorem 2.2.
+We only need to show Lipschitz continuity (2.10) holds, given the smoothness of $\mu_{i}$ and $\nu_{i}$.
+[^3] In most cases of interest, when $\mathbb{E}(X|u)$ is smooth, $\mathbb{E}(X^{2}|u)$ is smooth as well.
+Intuitively, the requirement that $\mathbb{E}(X^{2}|u)$ be smooth is the price we pay for not assuming boundedness of $X$.
 The proof is proceeded by induction.
 The base case holds, i.e.
-$|E(X_{i+1}|u_{i})-E(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
+$|\mathbb{E}(X_{i+1}|u_{i})-\mathbb{E}(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
 In the following, we investigate stochastic process of $u_{n}$, $n>i$. With $u_{i}$ is given, which may be treated as a deterministic variable, we write $\mathbb{E}$ for $\mathbb{E}_{i}$ .
@@
 $\mathrm{Var}_{n+1}=\mathbb{E}\big(u+x_{i+1}\cdots+x_{n}-(u+\bar{x}_{i+1}+\cdots+\bar{x}_{n})+(x_{n+1}-\bar{x}_{n+1})\big)^{2},$
-the following recurrence equation about $\mathrm{Var}_{n+1}$ holds.
+the following recurrence equation about $\mathrm{Var}_{n+1}$ holds,
 $\mathrm{Var}_{n+1}=\mathrm{Var}_{n}+2\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]+\frac{1}{N^{2}}\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ (2.14)
@@
 where the term of $\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$ is neglected.
-Now let’s estimate $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$. From (2.13) and (2.12), we obtain
+Now let’s estimate $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$. From (2.13) and (2.12), we have
 $\mathbb{E}(X_{n+1}^{2})-\overline{X}_{n+1}^{2}=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-[\mu_{{n}}(\bar{u}_{n})+\frac{1}{2}\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}]^{2}$
@@
 (higher order terms are neglected)
-Given above, it is easy to check that, in the region where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$.
+Given above, it is easy to check that, in the domain where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$.
 Further, by induction, $\frac{d\mathrm{Var}_{n+1}}{du}$ exists and its order is $O(1/N)$.
@@
 The average $(\cdot)$ over the whole population is
-$\mathbb{E}^{\!{}^{(M)}}[\,\cdot\,]=\frac{\displaystyle\int(\,\cdot\,)\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}{\displaystyle\int\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}$
+$\mathbb{E}^{\!{}^{(M)}}[\,\cdot\,]=\frac{\displaystyle\int(\,\cdot\,)\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}{\displaystyle\int\prod_{j=1}^{M}m_{j}\ P(dX_{1}\cdots dX_{\!{}_{M}})}$ (3.2)
 where $m_{i}=m_{i}(u_{i-1}+X_{i}/N)$. This is dependent of the future generations, whereas statistics in RW is independent of future. Therefore any statistical measure in BRW is generation dependent.
@@
 for a constant $\delta>0$ , $i=1,2,\cdots,$ then
-$Pr[\ |S_{n}-\mathbb{E}^{{(M)}}(S_{n})\mid\geq\lambda]\leq 2\exp{\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)},\quad\text{$n\leq M$ }$ (3.7)
+$Pr[\ |S_{n}-\mathbb{E}(S_{n})|\geq\lambda]\leq 2\exp{\left(-\frac{\delta^{2}}{4K^{2}}\frac{\lambda^{2}}{n}\right)},\quad\text{$n\leq M$ }$ (3.8)
+where $\mathbb{E}(S_{n})$ is given by (3.2).
 Remark.
@@
 Theorem 3.1.
-If $E(e^{\delta A|X_{i}|}\mid u_{i-1})\leq K$, then in the area where $m_{i}(u)$, $\mu_{i}(u)$ and $\nu_{i}(u)$ are smooth, the concentration inequality holds.
+If $E(e^{\delta A|X_{i}|}\mid u_{i-1})\leq K$, then in the domain where $m_{i}(u)$, $\mu_{i}(u)$ and $\nu_{i}(u)$ are smooth, the concentration inequality holds.
 Proof.
@@
 Let $f_{i}^{(0)}(u_{i})$ denote $m(u_{i})$ and write
-$\overline{X}_{n+1}^{(k)}=\mathbb{E}_{n}^{(k)}(X_{n+1}),\quad f_{n+1}^{(k-1)}(u_{n+1})=\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right];$
+$\overline{X}_{n+1}^{(k)}\equiv\mathbb{E}_{n}^{(k)}(X_{n+1}),\quad f_{n+1}^{(k-1)}(u_{n+1})\equiv\mathbb{E}_{n+1}^{(k-1)}\left[\,m(u_{n+k})\,\right];$
 thus (3.13) can be written as
@@
 Observe that the term $\frac{\overline{X}_{n+1}^{(k)}}{N}{f_{n+1}^{(k-1)}}^{\prime}(u_{n})$ involves no second order derivatives of $f_{n+2}^{(k-2)}$ which is absorbed into the $O(N^{-2})$ term.
-Thus, we have proved that for any $k$, $\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]$ is smooth in $u_{n}$.
+Thus, we have proved that for any $k$, $\mathbb{E}_{n}^{(k)}\left[\,m(u_{n+k})\,\right]$ is smooth in $u_{n}$, where
 $\overline{X}_{n+1}^{(k)}:=\mathbb{E}_{n}^{(k)}[\,X_{n+1}\,\mid u_{n}]=\frac{\displaystyle\int X_{n+1}\,m_{n+1}m_{n+2}\cdots m_{k}\,P(dX_{n+1}...dX_{k}\mid u_{n})}{\displaystyle\int m_{n+1}m_{n+2}\cdots m_{k}\,P(dX_{n+1}\cdots dX_{k}\mid u_{n})}$ (3.16)
+Hence, from (3.15),
 $\frac{\mathbb{E}_{n}^{(k+1)}\left[\,m(u_{n+1+k})\,\right]}{\mathbb{E}_{n+1}^{(k)}\big[\,m(u_{n+1+k})|X_{n+1}=0\,\big]}$ $=\frac{\displaystyle f_{n+1}^{(k)}\!\left(u_{n}\right)+\frac{\overline{X}_{n+1}^{(k+1)}(u_{n})}{N}{f_{n+1}^{(k)}}^{\prime}\!\left(u_{n}\right)+O\!\left(N^{-2}\right)}{f_{n+1}^{(k)}\left(u_{n}+0\right)}$
@@
 In the situation of bounded $X_{i}$, a particle always stays around the mean trajectory as long as its birthplace $u_{i-1}$ is near that trajectory.
-In the case of unboundedness, however, a particle may move into regions near singularities (e.g., critical points in K–SAT) where $u_{i}$ is not smooth.
-In such regions, $\mathbb{E}(e^{\delta Xi})<K$ fails, among others.
+In the case of unboundedness, however, a particle may move into regions near singularities (e.g., critical points in K–SAT) where $u_{i}$ is not presumed smooth.
+In such regions, $\mathbb{E}(e^{\delta Xi})<K$ may fail, among others.
 The good news is, because concentration inequality holds in the neighborhood area of the mean trajectory—is at least $\epsilon$ far away from the singularity—the outliers account for a negligible proportion of the population and can be ignored.
@@
 Theorem 3.2.
-Let $\epsilon$ be small compared with $\delta$, $N^{1/2}\ll\lambda<\epsilon N$, and as before,
+Let $\epsilon$ be small compared with $\delta$, $N^{1/2}\lesssim\lambda<\epsilon N$, and as before,
 $S_{i}:=\sum_{t=1}^{i}X_{t},\quad\Delta S_{i}:=|S_{i}-\mathbb{E}S_{i}|,\quad u_{i}:=\frac{S_{i}}{N}$
 Define the “good neighborhood” around the mean path
-$G_{i-1}:=\Big(\bar{u}_{i-1}-\frac{\lambda}{N},\bar{u}_{i-1}+\frac{\lambda}{N}\Big),\quad\bar{u}_{i-1}:=\frac{\mathbb{E}S_{i-1}}{N}$
+$D_{i-1}:=\Big(\bar{u}_{i-1}-\frac{\lambda}{N},\bar{u}_{i-1}+\frac{\lambda}{N}\Big),\quad\bar{u}_{i-1}:=\frac{\mathbb{E}S_{i-1}}{N}$
 Assume as before
-$E\left[e^{\delta|X_{i}|}\,\big|\,u_{i-1}\in G_{i-1}\right]\leq K\quad\text{{(bounded MGF)}}$
+$\mathbb{E}\left[e^{\delta A|X_{i}|}\,\big|\,u_{i-1}\in D_{i-1}\right]\leq K\quad\textnormal{{(bounded MGF)}}$
 Then
@@
 $\iff\left|u_{i-1}-\bar{u}_{i-1}\right|\leq\frac{\lambda}{N}$
-$\iff u_{i-1}\in G_{i-1}$
-For the base case, we have $u_{0}\in G_{0}$ where
-$G_{0}:=\Big(u_{0}-\frac{\lambda}{N},u_{0}+\frac{\lambda}{N}\Big);$
-bounded MGF is satisfied for $i=1$. Thus concentration inequality holds,
+$\iff u_{i-1}\in D_{i-1}$
+For the base case, we have $u_{0}\in D_{0}$ where
+$D_{0}:=\Big(u_{0}-\frac{\lambda}{N},u_{0}+\frac{\lambda}{N}\Big);$
+bounded MGF is satisfied for $i=1$. Thus the concentration inequality holds as
 $\Pr\Big(|\,S_{1}-\mathbb{E}[S_{1}]\,|\leq\lambda\Big)\geq\Big(1-2e^{-{c\lambda^{2}}/1}\Big)$
@@
 Then, given
-$u_{1}\in G_{1},u_{2}\in G_{2},...,u_{n-1}\in G_{n-1}$
+$u_{1}\in D_{1},u_{2}\in D_{2},...,u_{n-1}\in D_{n-1}$
 the standard concentration inequality can be established for $S_{n}$ by the same argument as in Lemma 2.3 and Theorem 2.2 :
-$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\geq\lambda\,\Big|\,u_{1}\in G_{1},...,u_{n-1}\in G_{n-1}\Big)\leq 2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
+$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\geq\lambda\,\Big|\,u_{1}\in D_{1},...,u_{n-1}\in D_{n-1}\Big)\leq 2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
 so
-$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\leq\lambda\,\Big|\,u_{1}\in G_{1},...,u_{n-1}\in G_{n-1}\Big)\geq 1-2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
+$\Pr\Big(\big|\,S_{n}-\mathbb{E}S_{n}\,\big|\leq\lambda\,\Big|\,u_{1}\in D_{1},...,u_{n-1}\in D_{n-1}\Big)\geq 1-2\exp\Big(-c\frac{\lambda^{2}}{n}\Big)$
 Thus

### v6

@@
 However, to our knowledge, no existing work eliminates both the martingale and boundedness conditions while achieving sub-Gaussian concentration; related work by Wormald (1995) is discussed later.
-The first major contribution of this paper is to do so.
-Informally, the small per-step effect in a random process can be controlled by the Lipschitz continuity of the mean increment with respect to the history, not necessarily by “almost independence”.
+The first major contribution of this paper is to do so; informally, under the condition of “finite” increment, the small per-step effect in a random process can be controlled by the Lipschitz continuity of the mean increment with respect to the history, not necessarily by “almost independence”.
 1.2 Notation and contributions
@@
 To connect the calculus notation with the underlying stochastic process, we define
-$\frac{\partial y}{\partial t}:=\frac{\mathbb{E}_{i}(y_{t})-\mathbb{E}_{i}(y_{t-1})}{1/N}=\mathbb{E}_{i}(Y_{t})-\mathbb{E}_{i}(Y_{t-1})=E_{i}(X_{t})$ (1.2)
+$\frac{\partial y}{\partial t}:=\frac{\mathbb{E}_{i}(y_{t})-\mathbb{E}_{i}(y_{t-1})}{1/N}=\mathbb{E}_{i}(Y_{t})-\mathbb{E}_{i}(Y_{t-1})=E_{i}(X_{t})\;(\equiv\mathbb{E}(X_{t}\mid u_{i}))$ (1.2)
 where $Y_{t}=X_{1}+X_{2}+...+X_{t}$, and
@@
 $\frac{\partial^{2}y}{\partial u_{i}\,\partial t}=0$ (1.4)
-While, traditionally, concentration inequalities require the condition (1.4), one of the primary contributions of this work is the relaxation of this classical constraint to the condition:
+While classical concentration inequalities require the condition (1.4), one of the primary contributions of this work is the relaxation of this classical constraint to the condition:
 $\left|\frac{\partial^{2}y}{\partial u_{i}\,\partial t}\right|\leq L,\quad i=1,2,...,n$
@@
 as adopted by textbooks. Sub-gaussian condition is a key ingredient in the proof of concentration inequalities, as this ultimately leads to a bound of the form
-$\exp({\sim nt^{2}-t\lambda})$
-which, given $\lambda=o(n)$, is exponentially small, $\exp(-n^{p})$ with $0<p<1$, under the choice $t\sim n^{(p-1)/2}$.
+$\exp({\sim nt^{2}-t\lambda}),\quad$
+for $\lambda\sim n^{(1+p)/2}$ with $0<p<1$, which can be made exponentially small—of order $\exp(-\Theta(n^{p}))$—by choosing sufficiently small $t$.
+Intuitively, provided $\lambda$ is not too small, the exponent can be made exponentially small by choosing small $t$, as $t^{2}$ is of higher order than $t$.
 Proof (A non-sharp version of Azuma’s inequality).
@@
 $\leq\cdots\leq\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}\quad\mbox{(cf. Azuma's $\exp{\left(\displaystyle-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$)}$
-The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ (owing to [34] and [43]) without compromising strength of results; from the practical point of view, the difference between the constant factors in the exponent, i.e., $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ versus $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$, is immaterial – both are square-exponential decay.
+The emphasis in this exposition is to show that the condition of $X_{i}<\infty$ can be relaxed to $E(e^{|X_{i}|})<\infty$ without compromising strength of results; from the practical point of view, the difference between the constant factors in the exponent, i.e., $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{2e^{c}}\right)}$ versus $\exp{\left(-\frac{\lambda^{2}}{2n}\cdot\frac{1}{c^{2}}\right)}$, is immaterial – both are square-exponential decay.
 In addition, it is shown that sub-Gaussian tail bounds can be obtained without “nifty” techniques.
-Next, we show what weaker conditions can ensure concentration inequalities, beyond independence, martingale differences, and boundedness. To this end, define Doob’s (or McDiarmid’s) martingale
-$d_{i}$ $=\mathbb{E}_{i}(X_{i}+\cdots+X_{n})-\mathbb{E}_{i-1}(X_{i}+\cdots+X_{n})$ (2.4)
+Next, we identify weaker conditions that ensure concentration inequalities, beyond independence, martingale differences, and boundedness. To this end, define Doob’s (or McDiarmid’s) martingale,
+$\Delta_{i}$ $=\mathbb{E}_{i}(X_{i}+\cdots+X_{n})-\mathbb{E}_{i-1}(X_{i}+\cdots+X_{n})$
 It is easy to check
-$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}d_{i}\,\,,\quad E_{i-1}(d_{i})=0$
+$S_{n}-\mathbb{E}S_{n}=\sum_{i=1}^{n}\Delta_{i}\,\,,\quad E_{i-1}(\Delta_{i})=0$
 Note
-$d_{i}$ $=X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)$ (2.5)
-Given $u_{i-1}$, $E_{i}(X_{l})$ is a function of $X_{i}$. $E_{i-1}(X_{l})=E_{X_{i}}\{E_{i}(X_{l})\}$ where $E_{X_{i}}(\cdot)$ takes expectation with respect to $X_{i}$. We have
+$\Delta_{i}$ $=X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)$ (2.4)
+Given $u_{i-1}$, $E_{i}(X_{l})$ is a function of $X_{i}$. $E_{i-1}(X_{l})=E_{X_{i}}\Big[E_{i}(X_{l})\Big]$ where $E_{X_{i}}(\cdot)$ takes expectation with respect to $X_{i}$. We have
 Lemma 2.2.
@@
 $|\mathbb{E}_{i}(X_{\ell})-\mathbb{E}_{i-1}(X_{\ell})|\leq L\frac{|X_{i}|}{N}+L\frac{\mathbb{E}_{i-1}(\,|X_{i}|\,)}{N}$ (2.6)
-Note, $\mathbb{E}_{i}(X_{\ell})\equiv\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$. Given $u_{i-1}$, this is a function of $X_{i}$.
+Note, $\mathbb{E}_{i}(X_{\ell})\equiv\mathbb{E}(X_{\ell}\mid X_{i},u_{i-1})$, which is a function of $X_{i}$ given $u_{i-1}$.
 Proof (Lemma 2.2).
@@
 $\mathbb{E}_{X_{i}^{\prime}}(\,\cdot\,)\equiv\mathbb{E}_{X_{i}^{\prime}}(\,\cdot\mid u_{i-1}),\quad X_{i}^{\prime}\sim\mathcal{L}(X_{i}\mid u_{i-1}).$
-With this notation, we have
+With this notation and the tower property of conditional expectation, we have,
 $\bigl|\mathbb{E}_{i}\left(X_{\ell}\right)-\mathbb{E}_{i-1}\left(X_{\ell}\right)\bigr|$ $=\Bigl|\,\mathbb{E}_{i-1}\left(X_{\ell}\mid X_{i}\right)-\mathbb{E}_{X_{i}^{\prime}}\!\left[\,\mathbb{E}_{i-1}\!\left(\,X_{\ell}\mid X_{i}^{\prime}\right)\right]\Bigl|$
@@
 Remarks.
-  • Condition (2.8) in Lemma 2.3 is equivalent to the following Lipschitz continuity
+  • Condition (2.7) in the Lemma 2.3 is equivalent to the following Lipschitz continuity
     $\left|\mathbb{E}(X_{\ell}\mid u_{i})-\mathbb{E}(X_{\ell}\mid u_{i}^{\prime})\right|\leq L\left|u_{i}-u_{i}^{\prime}\right|\quad\mbox{for $i<l$}$ (2.9)
@@
     $u_{i}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}),\quad u_{i}^{\prime}=\displaystyle\frac{1}{N}(X_{1}+...+X_{i-1}+X_{i}^{\prime}).$
+    Informally, the condition (2.7) asserts that the drift is Lipschitz continuous with respect to history, including the initial state $S_{0}$.
+We will later show that it is true if the drift is smooth.
   • For $n=O(N)$, the L-Lipschitz condtion may imply
@@
   • Independence and martingale difference assumptions both imply this L-Lipschitz condition (2.7). That is, the Lipschitz condition is weaker than those assumptions.
-  • Poisson, Binomial, and Geometric distributions, for example, all satisfy the finite exponential moment condition, as does the Gaussian distribution.
+  • Poisson(d)(d), Binomial(N,d/N)(N,d/N), and the variable occurrence distribution in SAT/COL models, for example, all satisfy the finite exponential moment condition, as does the Gaussian distribution.
 Proof of Lemma 2.3.
-By definition (2.5),
-$\mathbb{E}_{i-1}\left[e^{\delta|d_{i}|}\right]=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
-$\leq\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|\mathbb{E}_{i-1}(X_{i})|+\delta\sum_{\ell=i+1}^{n}\big|\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big|\Big)\Big]$
+By definition (2.4), we have
+$\mathbb{E}_{i-1}\left[e^{\delta|\Delta_{i}|}\right]=\mathbb{E}_{i-1}\exp\left(\delta\,\Big|X_{i}-\mathbb{E}_{i-1}X_{i}+\sum_{\ell=i+1}^{n}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)\,\Big|\right)$
+$\leq\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|+\delta|\mathbb{E}_{i-1}(X_{i})|+\delta\sum_{\ell=i+1}^{n}\left|\,\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\,\right|\right)$
+$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|+\delta\sum_{\ell=i+1}^{n}L\frac{|X_{i}|}{N}+\delta\sum_{\ell=i+1}^{n}L\frac{\mathbb{E}_{i-1}|X_{i}|}{N}\right)$
+(Lemma 2.2)
+$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|+\delta|X_{i}|L\frac{M}{N}+\delta\,\mathbb{E}_{i-1}(\,|X_{i}|\,)\cdot L\frac{M}{N}\right)$
+$=\exp\left(\delta\mathbb{E}_{i-1}|X_{i}|\,+\,\delta L\frac{M}{N}\mathbb{E}_{i-1}|X_{i}|\right)\cdot\mathbb{E}_{i-1}\exp\left(\delta|X_{i}|\Big(1+L\frac{M}{N}\Big)\right)$
+$=\exp\big(\delta A\mathbb{E}_{i-1}|X_{i}|\big)\cdot\mathbb{E}_{i-1}\left[e^{\delta A|X_{i}|}\right]\leq K^{2}$
+(Jensen’s inequality and the hypothesis)
+Applying the exponential Markov inequality $\Pr[X\geq\lambda]\leq e^{-t\lambda}E(e^{tX})$ and noting
+$\mathbb{E}_{i-1}(\Delta_{i})=0$
+we obtain
+$\Pr\big(S_{n}-\mathbb{E}S_{n}\geq\lambda\,\big)$ $\leq e^{-t\lambda}\,\mathbb{E}\!\left[e^{\,t(S_{n}-\mathbb{E}S_{n})}\right]$
+$=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(\Delta_{1}+\Delta_{2}+\cdots+\Delta_{n})}\right]$
+$=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(\Delta_{1}+\cdots+\Delta_{n-1})}\;\mathbb{E}_{n-1}\left(e^{t\Delta_{n}}\right)\right]$
 […]
-$\leq e^{\mathbb{E}_{i-1}(\delta|X_{i}|)}\,\mathbb{E}_{i-1}\Big[\exp\Big(\delta|X_{i}|+\delta|X_{i}|L\frac{M}{N}+\delta\,\mathbb{E}_{i-1}(\,|X_{i}|\,)\cdot L\frac{M}{N}\Big)\Big]$
-$=\exp\left(\mathbb{E}_{i-1}\delta|X_{i}|\,+\,\delta L\frac{M}{N}\mathbb{E}_{i-1}|X_{i}|\right)\cdot\mathbb{E}_{i-1}\left[\exp\left(\delta|X_{i}|\Big(1+L\frac{M}{N}\Big)\right)\right]$
-$=\exp\big(\mathbb{E}_{i-1}\delta\cdot A|X_{i}|\big)\cdot\mathbb{E}_{i-1}(e^{\delta A|X_{i}|})\leq K^{2}~~~~\mbox{(Jensen's inequality and the hypothesis)}$
-Applying the exponential Markov inequality $\Pr[X\geq\lambda]\leq e^{-t\lambda}E(e^{tX})$ and noticing
-$\mathbb{E}_{i-1}(d_{i})=0$
-it follows
-$\Pr\big(S_{n}-\mathbb{E}S_{n}\geq\lambda\,\big)$
-$\leq e^{-t\lambda}\,\mathbb{E}\!\left[e^{\,t(S_{n}-\mathbb{E}S_{n})}\right]=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(d_{1}+d_{2}+\cdots+d_{n})}\right]=e^{-t\lambda}\,\mathbb{E}\!\left[e^{t(d_{1}+\cdots+d_{n-1})}\;\mathbb{E}_{n-1}\left(e^{td_{n}}\right)\right]$
-[…]
-$\leq e^{-t\lambda}\,e^{\,t^{2}K^{2}/\delta^{2}}\;\mathbb{E}\!\left[e^{t(d_{1}+\cdots+d_{n-2})}\;\mathbb{E}_{n-2}\!\left(e^{td_{n-1}}\right)\right]$
+$\leq e^{-t\lambda}\,e^{\,t^{2}K^{2}/\delta^{2}}\;\mathbb{E}\!\left[e^{t(\Delta_{1}+\cdots+\Delta_{n-2})}\;\mathbb{E}_{n-2}\!\left(e^{t\Delta_{n-1}}\right)\right]$
 $\quad\vdots$
@@
   • The condition of Lipschitz continuity (2.9) may be interpreted as “continuity” of the trajectory with respect to the history; i.e.
-small differences in initial position do not lead to vehement differences in the trajectory.
+small differences in initial position do not lead to drastic differences in the trajectory.
   • In the case of independence, $\mathbb{E}_{i}[X_{\ell}]=\mathbb{E}[X_{\ell}]$, and hence $\mathbb{E}_{i}[X_{\ell}]-\mathbb{E}_{i-1}[X_{\ell}]$ = 0.
@@
     $\sum_{\ell>i}\big(\mathbb{E}_{i}X_{\ell}-\mathbb{E}_{i-1}X_{\ell}\big)=0$
-    and therefore $d_{i}$ (see, (2.5)) is reduced to $X_{i}-\mathbb{E}X_{i}$ (independence) and to $X_{i}$ (martingale difference), respectively.
+    and therefore $\Delta_{i}$ (see, (2.4)) reduces to $X_{i}-\mathbb{E}X_{i}$ (independence) and to $X_{i}$ (martingale difference), respectively.
 In these cases, the condition of bounded difference or the existence of exponential moment alone suffices to imply the concentration inequalities.
@@
   • Two special cases of (2.9) are independence where $\frac{\partial y}{\partial t}$ is constant and martingale difference where $\frac{\partial y}{\partial t}=0$.
-The following theorem provides easily verifiable conditions for concentration inequality in general cases where
+The following theorem provides readily verifiable conditions for concentration inequality in general settings where
 $\frac{\partial}{\partial u_{i}}\left(\frac{\partial y}{\partial t}\right)\neq 0$
-i.e. $\frac{\partial y}{\partial t}$ is not constant with $u_{i}$ varying.
+i.e. $\frac{\partial y}{\partial t}$ is not necessarily constant with $u_{i}$ varying.
 $\mathbb{E}(X_{i+1}|u_{i})$ is a function of $u_{i}$. Let $\mu_{i}(u_{i})$ denote $E(X_{i+1}|u_{i})$ and $\nu_{i}(u_{i})$ denote $E(X_{i+1}^{2}|u_{i})$;
@@
 Let $\mu_{i}^{\prime}$ denote $\frac{d\mu_{i}(u_{i})}{du_{i}}$, $\nu_{i}^{\prime}$ denote $\frac{d\nu_{i}(u_{i})}{du_{i}}$, $\mu_{i}^{\prime\prime}$ the second order derivative, and so on.
-We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the domain in question lies sufficiently away from singularities).
+We shall use term “smooth function” to refer that (a) the first several order derivatives (with respective to $u_{i}$, briefly $u$) exist and (b) they are uniformly bounded (in other words, the domain in question is bounded away from the singularities).
 Theorem 2.2.
 If $\mathbb{E}\big(e^{\delta A|X_{i}|}\mid u_{i-1}\big)\leq K$ (existence of an exponential moment), then in the domain where $\mu_{i}$ and $\nu_{i}$ are smooth, the concentration inequality holds.
-In the notation of partial derivatives, the condition in Theorem 2.2 states that
+In terms of partial derivatives, the condition in Theorem 2.2 states that
 $\frac{\partial\mu(t,u)}{\partial u},\quad\frac{\partial^{2}\mu(t,u)}{\partial u^{2}},\quad\frac{\partial\nu(t,u)}{\partial u},\quad\cdots$
-are uniformly bounded in the domain of smoothness (i.e., away from singularities), in addition to the exponential moment condition.
+are uniformly bounded on the domain (i.e., away from singularities), in addition to the exponential moment condition.
 These conditions are practically easy to verify because we can examine $\bar{u}$, the mean of $u$, when in particular it satisfies a partial differential equation (PDE).
@@
 Proof of Theorem 2.2.
-We only need to show Lipschitz continuity (2.10) holds, given the smoothness of $\mu_{i}$ and $\nu_{i}$.
-[^3] In most cases of interest, when $\mathbb{E}(X|u)$ is smooth, $\mathbb{E}(X^{2}|u)$ is smooth as well.
-Intuitively, the requirement that $\mathbb{E}(X^{2}|u)$ be smooth is the price we pay for not assuming boundedness of $X$.
-The proof is proceeded by induction.
-The base case holds, i.e.
-$|\mathbb{E}(X_{i+1}|u_{i})-\mathbb{E}(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
+We only need to show Lipschitz continuity (2.9) holds, given the smoothness of $\mu_{i}$ and $\nu_{i}$$\dagger$$\dagger\daggerIn$ most cases of interest, when $\mathbb{E}(X|u)$ is smooth, $\mathbb{E}(X^{2}|u)$ is smooth as well.
+Intuitively, regularity of $\mathbb{E}(X^{2}|u)$ is the price paid for dispensing with a boundedness assumption on $X$.
+The proof is proceeded by induction. The base case holds, i.e. $|\mathbb{E}(X_{i+1}|u_{i})-\mathbb{E}(X_{i+1}|u_{i}^{\prime})|\leq L|u_{i}-u_{i}^{\prime}|$, because of the smoothness of $\mu_{i}$.
 In the following, we investigate stochastic process of $u_{n}$, $n>i$. With $u_{i}$ is given, which may be treated as a deterministic variable, we write $\mathbb{E}$ for $\mathbb{E}_{i}$ .
@@
 and therefore we have
-$\mathbb{E}(X_{n+1})=\mu_{n}(\bar{u}_{n})~+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\frac{C}{N^{3/2}}$ (2.12)
+$\mathbb{E}(X_{n+1})=\mu_{n}(\bar{u}_{n})~+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\sim\frac{1}{N^{3/2}}$ (2.11)
 where $\mathrm{Var}_{n}$ denotes $\mathbb{E}(u_{n}-\bar{u}_{n})^{2}$. Similarly, for $\nu$ we have
-$\mathbb{E}(X_{n+1}^{2})=$ $\nu_{n}(\bar{u}_{n})+\frac{1}{2}\nu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+~\frac{C}{N^{3/2}}$ (2.13)
+$\mathbb{E}(X_{n+1}^{2})=$ $\nu_{n}(\bar{u}_{n})+\frac{1}{2}\nu_{n}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\sim\frac{1}{N^{3/2}}$ (2.12)
 From
 $\mathrm{Var}_{n+1}=\mathbb{E}\big(u+x_{i+1}\cdots+x_{n}-(u+\bar{x}_{i+1}+\cdots+\bar{x}_{n})+(x_{n+1}-\bar{x}_{n+1})\big)^{2},$
-the following recurrence equation about $\mathrm{Var}_{n+1}$ holds,
+the following recurrence equation about $\mathrm{Var}_{n}$ holds,
 $\mathrm{Var}_{n+1}=\mathrm{Var}_{n}+2\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]+\frac{1}{N^{2}}\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ (2.13)
-We first estimate the second term in the right hand-side of (2.14).
-$\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mathbb{E}[X_{n+1}(u_{n}-\bar{u}_{n})]-\bar{x}_{n+1}(\bar{u}_{n}-\bar{u}_{n})$
+Applying the law of total expectation with $\mathbb{E}[X_{n+1}\,|\,u_{n}]=\mu_{n}(u_{n})$, the cross term in the right hand-side of (2.13) is
+$\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mathbb{E}\left[\,X_{n+1}(u_{n}-\bar{u}_{n})\,\right]-\mathbb{E}\left[\,\bar{x}_{n+1}(u_{n}-\bar{u}_{n})\,\right]$
 $=$ $\frac{1}{N}\mathbb{E}[\mu_{n}(u_{n})(u_{n}-\bar{u}_{n})]$
-Again applying Taylor expansion to $\mu_{n}(u_{n})$, i.e.
+Now from the Taylor expansion of $\mu_{n}(u_{n})$ around $\bar{u}$, i.e.
 $\mu_{n}(u_{n})=\mu_{n}(\bar{u}_{n})+\mu_{n}^{\prime}(\bar{u}_{n})(u_{n}-\bar{u}_{n})+\frac{1}{2}\mu_{n}^{\prime\prime}(\bar{u}_{n})(u_{n}-\bar{u}_{n})^{2}+\sim(u_{n}-\bar{u}_{n})^{3}$
-we have, with the term of $\sim\mathbb{E}(u_{n}-\bar{u}_{n})^{4}$ neglected,
+neglecting higher-order terms, $\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$ and terms of order $\mathbb{E}(u_{n}-\bar{u}_{n})^{4}$, we have,
 $\mathbb{E}\left[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})\right]=$ $\frac{1}{N}\mu_{n}(\bar{u}_{n})\mathbb{E}(u_{n}-\bar{u}_{n})+\frac{1}{N}\mu_{n}^{\prime}(\bar{u}_{n})\mathrm{Var}_{n}+\frac{1}{2N}\mu_{n}^{\prime\prime}(\bar{u}_{n})\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$
 $=$ $\frac{1}{N}\mu_{n}^{\prime}(\bar{u}_{n})\mathrm{Var}_{n}$
-where the term of $\mathbb{E}(u_{n}-\bar{u}_{n})^{3}$ is neglected.
+where we have used $\mathbb{E}(u_{n}-\bar{u}_{n})=0$ and omitted the cubic terms as negligible.
 Now let’s estimate $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$. From (2.12) and (2.11), we have
-$\mathbb{E}(X_{n+1}^{2})-\overline{X}_{n+1}^{2}=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-[\mu_{{n}}(\bar{u}_{n})+\frac{1}{2}\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}]^{2}$
+$\mathbb{E}(X_{n+1}^{2})-\overline{X}_{n+1}^{2}=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-\left[\mu_{{n}}(\bar{u}_{n})+\frac{1}{2}\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}\right]^{2}$
 $=$ $\nu_{{n}}(\bar{u}_{n})+\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}-\mu_{{n}}^{2}(\bar{u}_{n})-\mu_{{n}}(\bar{u}_{n})\mu_{{n}}^{\prime\prime}(\bar{u}_{n})\mathrm{Var}_{n}$
 $=$ $\nu_{{n}}(\bar{u}_{n})-\mu_{{n}}^{2}(\bar{u}_{n})+\left(\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})-\mu_{n}(\bar{u}_{n})\mu_{n}^{\prime\prime}(\bar{u}_{n})\right)\mathrm{Var}_{n}$
-For illustration, at $i+1$
+For illustration, at $i+1$,
 $\mathbb{E}(X_{i+1}-\overline{X}_{i+1})^{2}=\mathbb{E}(X_{i+1}^{2})-\mu_{i}^{2}=\nu_{i}(u_{i})-\mu_{i}^{2}(u_{i})=\mathrm{Var}(X_{i+1})$
 where $\mathrm{Var}(X_{i})=0$ because $u_{i}$ is treated as a deterministic variable.
-By substituting $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ and $\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]$ into (2.14), we have a recurrence relation for $n\geq i$,
+For final variance recurrence, substituting $\mathbb{E}(X_{n+1}-\overline{X}_{n+1})^{2}$ and $\mathbb{E}[(x_{n+1}-\bar{x}_{n+1})(u_{n}-\bar{u}_{n})]$ back into (2.13), we have the recurrence relation for $n\geq i$,
 $\mathrm{Var}_{n+1}=$ $\mathrm{Var}_{n}+\frac{2}{N}\mu_{n}^{\prime}(\bar{u}_{n})\mathrm{Var}_{n}~+\frac{1}{N^{2}}\left[\nu_{{n}}(\bar{u}_{n})-\mu_{{n}}^{2}(\bar{u}_{n})+\left(\frac{1}{2}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})-\mu_{n}(\bar{u}_{n})\mu_{n}^{\prime\prime}(\bar{u}_{n})\right)\mathrm{Var}(u_{n})\right]$
 $=$ $\mathrm{Var}_{n}+\frac{\mathrm{Var}_{n}}{N}\left(2\mu_{n}^{\prime}(\bar{u}_{n})~+\frac{1}{2N}\nu_{{n}}^{\prime\prime}(\bar{u}_{n})-\frac{1}{N}\mu_{n}(\bar{u}_{n})\mu_{n}^{\prime\prime}(\bar{u}_{n})\right)+\frac{1}{N^{2}}\left(\nu_{n}(\bar{u}_{n})-\mu_{n}^{2}(\bar{u}_{n})\right)$
-and $\mathrm{Var}_{i}=0$. With the argument $\bar{u}_{n}$” omitted for brevity:
+and the base case is $\mathrm{Var}_{i}=0$. With the argument $\bar{u}_{n}$” omitted for brevity:
 $\mathrm{Var}_{n+1}=$ $\mathrm{Var}_{n}\left[1+\frac{1}{N}\left(2\mu_{n}^{\prime}+\frac{1}{2N}\nu_{n}^{\prime\prime}-\frac{1}{N}\mu_{n}\mu_{n}^{\prime\prime}\right)\right]+\frac{1}{N^{2}}(\nu_{n}-\mu_{n}^{2})$
-Define
+We define the coefficients
 $a_{n}=\displaystyle 1+\frac{1}{N}\left(2\mu_{n}^{\prime}+\frac{1}{2N}\nu_{n}^{\prime\prime}-\frac{1}{N}\mu_{n}\mu_{n}^{\prime\prime}\right),\qquad b_{n}=\frac{\nu_{n}-\mu_{n}^{2}}{N^{2}}$
-Then, the above recursion for $\mathrm{Var}$ gives
-$\mathrm{Var}_{n+1}=$ $\sum_{j=i+1}^{n-1}b_{j}\cdot a_{j+1}...a_{n}+b_{n}=\sum_{j=i+1}^{n-1}b_{j}\cdot a_{j+1}...a_{n}+\frac{1}{N^{2}}(\nu_{n}-\mu_{n}^{2})$
-$=$ $\frac{1}{N^{2}}\left[\sum_{j=i+1}^{n-1}(\nu_{j}-\mu_{j}^{2})\prod_{k=j+1}^{n}\left(1+\frac{1}{N}\Big(2\mu_{k}^{\prime}+\frac{1}{2N}\nu_{k}^{\prime\prime}-\frac{1}{N}\mu_{k}\mu_{k}^{\prime\prime}\Big)\right)+\nu_{n}-\mu_{n}^{2}\right]$
-$=$ $\frac{1}{N^{2}}\left[\sum_{j=i+1}^{n-1}\left(\nu_{j}-\mu_{j}^{2}\right)\prod_{k=j+1}^{n}\left(1+\frac{2\mu_{k}^{\prime}}{N}\right)\right]$
-(higher order terms are neglected)
-Given above, it is easy to check that, in the domain where $\mu_{j}^{\prime}$, $\mu_{j}^{\prime\prime}$ and $\nu_{j}^{\prime\prime}$ are bounded, j=1,2nj=1,2...n, $\mathrm{Var}_{n+1}$ is bounded by a constant of $O(1/N)$.
-Further, by induction, $\frac{d\mathrm{Var}_{n+1}}{du}$ exists and its order is $O(1/N)$.
-From the base that $\Delta u_{i}=u_{i}-u_{i}^{\prime}$, and the inductive hypothesis that $\Delta\bar{u}_{n-1}$ is small, it follows from (2.12) that
-$\Delta\mathbb{E}(X_{n})$ $=\Delta\mu_{n-1}~+~\Delta\frac{1}{2}\mu_{n-1}^{\prime\prime}\mathrm{Var}_{n-1}~+~\frac{C}{N^{3/2}}$
+Then, the variance recurrence relation for $n\geq i$ takes the compact form
+$\mathrm{Var}_{n+1}=a_{n}\cdot\mathrm{Var}_{n}+b_{n}$
+It follows, for $n\geq i+1$
+$\mathrm{Var}_{n}=$ $\mathrm{Var}_{i}\prod_{k=i}^{n-1}a_{k}+\sum_{j=i}^{n-1}\Bigg(b_{j}\prod_{k=j+1}^{n-1}a_{k}\Bigg)$
+$=$ $\frac{1}{N^{2}}\left[\sum_{j=i}^{n-1}(\nu_{j}-\mu_{j}^{2})\prod_{k=j+1}^{n-1}\left(1+\frac{1}{N}\Big(2\mu_{k}^{\prime}+\frac{1}{2N}\nu_{k}^{\prime\prime}-\frac{1}{N}\mu_{k}\mu_{k}^{\prime\prime}\Big)\right)\right]$
+$=$ $\frac{1}{N^{2}}\left[\sum_{j=i}^{n-1}\left(\nu_{j}-\mu_{j}^{2}\right)\prod_{k=j+1}^{n-1}\left(1+\frac{2\mu_{k}^{\prime}}{N}\right)\right]$
+where higher order terms in $N^{-1}$ have been neglected.
+[…]
+From the inductive hypothesis that $|\Delta\bar{u}_{n-1}|\leq L_{n-1}|u_{i}-u_{i}^{\prime}|=O(\Delta u_{i})$, so that $\Delta\bar{u}_{n-1}$ is small, it follows from (2.11) that, at step $n$,
+$\Delta\mathbb{E}(X_{n})=\Delta\mu_{n-1}~+~\Delta\frac{1}{2}\mu_{n-1}^{\prime\prime}\mathrm{Var}_{n-1}~+~\frac{C}{N^{3/2}}\,.$
+Expanding,
 $=\mu_{n-1}^{\prime}\Delta\bar{u}_{n-1}+\frac{1}{2}\mu_{n-1}^{\prime\prime\prime}\mathrm{Var}_{n-1}\Delta\bar{u}_{n-1}+\frac{1}{2}\mu_{n-1}^{\prime\prime}\frac{d\mathrm{Var}_{n-1}}{du}\Delta\bar{u}_{n-1}~+~O(\Delta\bar{u}_{n-1}^{2})$
@@
 $=\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O\left(N^{-1}\right)+O(\Delta\bar{u}_{n-1})\Big)$
-where the 3rd term in (2.12), of order $N^{-3/2}$, is omitted since it yields $\sim\displaystyle O(N^{-3/2})\Delta\bar{u}_{n-1}$ in the above equation. From $u_{n}=u_{n-1}+X_{n}/N$ and then
-$\bar{u}_{n}=\bar{u}_{n-1}+\frac{1}{N}E(X_{n})$
-we have
-$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta E(X_{n})=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O\left(N^{-1}\right)+O(\Delta\bar{u}_{n-1})\Big)$
+Here the third term in (2.11), of order $N^{-3/2}$, has been omitted since it contributes only $O(N^{-3/2})\Delta\bar{u}_{n-1}$ in the above expansion. From
+$u_{n}=u_{n-1}+\frac{X_{n}}{N},\qquad\bar{u}_{n}=\bar{u}_{n-1}+\frac{1}{N}\mathbb{E}(X_{n})$
+it follows
+$\Delta\bar{u}_{n}$ $=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\mathbb{E}(X_{n})$
+$=\Delta\bar{u}_{n-1}+\frac{1}{N}\Delta\bar{u}_{n-1}\Big(\mu_{n-1}^{\prime}+O(N^{-1})+O(\Delta\bar{u}_{n-1})\Big)$
 Thus,
-$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}\left(1+\frac{1}{N}\mu_{n-1}^{\prime}+o\Big(\frac{1}{N}\Big)\right)$
-Iterating on $n$ yields
-$|\,\mathbb{E}(X_{n}\mid u_{i})-\mathbb{E}(X_{n}\mid u_{i}^{\prime})\,|\leq L|u_{i}-u_{i}^{\prime}|\quad\text{(where $u_{i}-u_{i}^{\prime}=\Delta u_{i}$)}$
-This concludes the proof. ∎
-We see from the above proof that Lipschitz continuity of $\mathbb{E}(f(u_{n}))$ is ultimately down to the existence of the first several order derivatives of $f$.
+$\Delta\bar{u}_{n}=\Delta\bar{u}_{n-1}\left(1+\frac{1}{N}\mu_{n-1}^{\prime}+\frac{o(1)}{N}\right)$
+Iterating in $n$ yields
+$\left|\mathbb{E}(X_{n}\mid u_{i})-\mathbb{E}(X_{n}\mid u_{i}^{\prime})\right|\leq L_{n}\left|u_{i}-u_{i}^{\prime}\right|,\quad u_{i}-u_{i}^{\prime}=\Delta u_{i},$
+with $L_{n}=\Theta(L)$. This concludes the proof. ∎
+From the above proof, we see that Lipschitz continuity of $\mathbb{E}(f(u))$ with respect to the process history is down to the existence of the first few derivatives of $f$ and of the variance of $u$.
 3 Concentration inequalities in the context of BRW
@@
 $\big|\,\mathbb{E}^{\,\!{}^{(M)}}(X_{\ell}\mid u_{i})-\mathbb{E}^{\,\!{}^{(M)}}(X_{\ell}\mid u_{i}^{\prime})\,\big|\,\leq\,L|u_{i}-u_{i}^{\prime}|~~~~for~i<l$
-i.e. $E^{(M)}_{i}(X_{\ell})$ is Lipschitz continuous function of $u_{i}$.
+i.e. $\mathbb{E}^{(M)}_{i}(X_{\ell})$ is Lipschitz continuous function of $u_{i}$.
 Recall that in the proof of Theorem 2.2, smooth $\mu_{i}(u_{i})$ and $\nu_{i}(u_{i})$ are employed to prove “uniqueness of solution” (in RW “local” smoothness is “global” smoothness because it won’t change with generations—steps).
-In the case of BRW, we need smoothness of $\mu_{i}^{\,\!{}^{(M)}}(u_{i}):=E_{i}^{\,\!{}^{(M)}}(X_{i+1}|u_{i})$, $\nu_{i}^{\,\!{}_{(M)}}(u_{i}):=E_{i}^{\,\!{}^{(M)}}(X_{i+1}^{2}|u_{i})$, and of $\mathbb{E}_{i}^{(M)}\!\big(m(u_{i+M})\mid u_{i}\big)$.
+In the case of BRW, we need smoothness of $\mu_{i}^{\,\!{}^{(M)}}(u_{i}):=\mathbb{E}_{i}^{\,\!{}^{(M)}}(X_{i+1}|u_{i})$, $\nu_{i}^{\,\!{}_{(M)}}(u_{i}):=\mathbb{E}_{i}^{\,\!{}^{(M)}}(X_{i+1}^{2}|u_{i})$, and of $\mathbb{E}_{i}^{(M)}\!\big(m(u_{i+M})\mid u_{i}\big)$.
 To this end, we first transform the “local” smoothness (given as hypotheses in the theorem) into “global” smoothness; i.e.